### v1

# Current distribution and chirality in the regime of the quantized Hall effect

Abstract
The integer quantum Hall effect (QHE) is one of the most basic and, at the same time, astonishing phenomena in condensed matter physics. Despite its importance for metrology and fundamental physics, our microscopic understanding of the QHE is not yet complete. For decades, there has been a discussion about the current distribution inside a Hall bar for the case of a quantized Hall resistance. With the advent of quantum technology, the questions, if the current flows in compressible Landauer-Büttiker edge channels or in incompressible strips with varying geometry and whether the current flow is chiral or not, gain importance, because the quantum phase of each charge carrier depends on its exact pathway. In the present work, we compare multi-terminal current measurements in the regime of the QHE with model calculations in order to unravel the current density distribution in a Hall-bar device. We find, that the conjecture of chirality is justified for the plateaus of quantized Hall resistance, while in between plateaus the current density looses its chirality. The latter is described by the Drude model, which predicts a homogeneous current density for a sample with homogeneous mobility and carrier density.
I Introduction
Magnetic field is an axial vector and, therefore, breaks time-reversal symmetry, which, however, can be restored by the simultaneous reversal of magnetic field and time, $\vec{B}\to-\vec{B}$ and $t\to-t$. This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$. The axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ in the classical equation of motion $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$ of an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$. As an introduction, we discuss the dynamics of charge carriers considering a variety of boundary conditions.
The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a constant drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf. Fig. 1(a). The chirality of this trajectory is left-handed for the negatively charged electron (but would be right handed for a positive charge). Note that we define the chirality of a trajectory in analogy to the handedness of a screw thread.
Figure 1: Motion (expressed by drift velocity) of an electron confined to the $xy$-plane with in-plane electric and perpendicular magnetic fields, $\vec{E}$ and $\vec{B}$ (red arrows), respectively. (a) Ballistic limit (free electron). The spiral is the classical trajectory consisting of a constant drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity. (b) Diffusive limit. For $|\mu B|<1$ the spiral motion is damped. However, the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ and tilts it by the Hall angle $\varphi$. (c) Diffusive Hall bar in the steady state limit with edges parallel to the applied $\vec{E}$ (red arrow). The Hall voltage compensates the Lorentz force, $-e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$ (or $E_{y}=|\vec{F}_{\text{L}}|/e$), such that the drift velocity is $\vec{\upsilon}_{\text{d}}$ as for $\vec{B}=0$.
The Drude model describes magnetotransport in the classical limit of the diffusive regime. Its key parameter is the electron mobility $\mu$. For $\mu B<1$ (corresponding to a mean-free path $l_{\text{m}}<R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering. Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$ and applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$, where $\tau=m\mu/e$ is the momentum scattering time. In this diffusive limit and for $\vec{B}\perp\vec{E}$, the solution of the equation of motion introduced above can be written as
$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,.$ (1)
[For $\vec{E}\not\perp\vec{B}$ there would be a third summand, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$, inside the parantheses.] Clearly, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous. However for finite $B$, $\vec{j}$ is not parallel to $\vec{E}$, but bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$, cf. Fig. 1(b).
The ballistic limit discussed above is clearly beyond the scope of the Drude model. Indeed, for $\mu B\to\infty$, the friction forces in Eq. (1) become negligible compared to the Lorentz force and the drift velocity becomes independent of $\mu$. Nevertheless for $\mu B\to\infty$, the drift velocity takes the value $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, which is identical to that of a ballistic electron. This observation encourages us to compare measurements with the Drude model even for $\mu B\gg 1$.
In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, and include boundaries of the 2DES and electron scattering. Furthermore, we consider an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular, homogeneous magnetic field $\vec{B}=(0,0,B)$, corresponding to one of the most common set-ups for magnetotransport measurements. For simplicity, we restrict our general discussion to a homogeneous electron density and mobilty inside the 2DES (away from the edges).
To discuss the classical limit of the Hall effect, we include two sample edges at $y=0$ and $y=b$, such that the electrons are confined in the $y$-direction but move freely in the $x$-direction, cf. Fig. 1(c). In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons. Applying this condition to Eq. (1) we find $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$. The transversal field $E_{y}=\vec{\nabla}V_{\text{H}}$ is maintained by charge accumulation at the sample edges, which gives rise to the Hall voltage $V_{\text{H}}$. In the steady state, $-eE_{y}=-F_{\text{L}}$, i.e., the Lorentz force acting on a free electron is exactly canceled, and the electrons move as if no magnetic field was applied but only the longitudinal field $E_{x}$. For our homogeneous sample, the current density $j_{x}$ is a constant and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
In a confined sample, the axial character of the magnetic field manifests itself in the edge charges or the related Hall voltage (which break the time reversal symmetry). The latter gives rise to an increase of the two-terminal resistance of the 2DES according to $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}$ with the resistance $R_{0}$ of the 2DES at $B=0$.
As we increase $B$, the quantization of the density of states into Landau-levels results in the quantized Hall effect (QHE). Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance. The latter indicates scattering-free transport within the Hall bar.
To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of perfect one-dimensional (1D) edge channels, in which scattering is suppressed. The conductance of a spin-resolved 1D channel is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$. The filling factor $\nu$ describes the number of perfect 1D channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk. The perfect 1D-channels are assumed to follow the edges of the Hall bar, because they form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2]. To obtain a quantized Hall resistance, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization. The direction of the current flow in the edge channels is predetermined by the gradient of the confinement potential. This yields current flow in opposite directions at opposite edges. In that sense, in the LBP the QHE is associated with chiral current flow.
The LBP provides an intuitive explanation for the quantized Hall resistance and the vanishing longitudinal resistance of the QHE. While its merit is its simplicity, its predictions are limited due to the phenomenological character as well as its neglect of interactions and non-equilibrium effects. In particular, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [3, 4]. However, this would cause highly charged stripes between the edge channels corresponding to an unrealistically large Coulomb energy.
The screening theory [3, 4, 5, 6, 7, 8, 9] overcomes these limitations by taking into account the direct Coulomb interaction between the electrons. The Coulomb interaction causes a rearrangement of the free carriers such that the highly charged stripes are avoided. Instead, it leads to a segmentation of the Hall bar in compressible regions and incompressible strips (ICSs). Inside the ICSs, carriers cannot scatter and screening of the electric field is absent, because the chemical potential is in a local energy gap between filled and empty Landau levels; ICSs are accompanied by adjacent compressible regions, in which a Landau-level is aligned with the chemical potential, such that carriers can scatter. Inside the compressible regions, the electric field is then perfectly screened. The current flows where the electric field drops, i.e., inside the ICSs. In contrast to the 1D Landauer-Büttiker edge channels, the ICSs have a finite width exceeding the magnetic length [7, 8, 9]. Width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$. In particular, towards the end of each plateau with increasing magnetic field, the screening theory predicts a single extended incompressible region in the center of the Hall bar [7, 8, 9]. This implies that current can flow in the bulk of the Hall bar and not just in edge channels. The predicted bulk ICS near the high magnetic field end of the plateaus was experimentally corroborated in scanning gate spectroscopy experiments [10, 11, 12] and recently directly confirmed in transport experiments [13]. As we will discuss in Sec. IV below, the screening theory is consistent with chiral transport at the plateaus of the QHE, even if the current flows inside a single bulk ICS.
In the present article, we investigate the influence of a magnetic field on the current density distribution of a multiterminal junction. In Sec. II we illuminate the non-chiral character of the Drude model and derive a general coherent and chiral model. In Sec. III we compare our multi-terminal current measurements with the predictions of the models. For relatively small magnetic fields, we find that the classical limit of the Hall effect is described by the Drude model. For higher magnetic fields, we distinguish between the plateau regions of a quantized Hall resistance and the transition regions between plateaus. For the plateaus, coherent and chiral transport was already supported by Mach-Zehnder interferometer experiments measuring interference between two distinct current paths [14]. Our own results confirm this. For the magnetic field regions in between the plateaus our measurements are consistent with the predictions of the classical Drude model, which predicts non-chiral transport. In Sec. IV we explain, why the screening theory predicts chiral transport.
II model
Figure 2: Drude model: Distribution of current $I=I_{xx}+I_{xy}$ in a three-terminal device with homogeneous electron density and mobility in the diffusive and classical limit for $V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{2}=R_{3}$ and $\vec{B}$ pointing into the plane of the 2DES. Color intensities indicate the distribution of the absolute value of the current density into the three branches. Within each branch, the current density is homogeneous. For $V_{1}<0$, white arrows show the directions of the conventional currents and black bent arrows the direction and relative magnitudes of electron flow (for $V_{1}>0$ vice versa). (a) $I_{xx}$ displaying ohmic behavior, the same as for $B=0$: $I_{xx}^{(2)}=I_{xx}^{(3)}=-0.5I_{xx}^{(1)}$. (b) $I_{xy}$ for $R_{2}=R_{3}=0$ displaying non-ohmic behavior: $I_{xy}$ is bent by the Lorentz force: $I_{xy}^{(2)}=-I_{xy}^{(1)}$ and $I_{xy}^{(3)}=0$. (c) $I_{xy}$ for $R_{2}=R_{3}>0$: $I_{xy}$ is still bent by the Lorentz force but partly reflected at the resistive contacts: $I_{xy}^{(2)}>I_{xy}^{(3)}$ with $I_{xx}^{(2)}+I_{xx}^{(3)}=-I_{xx}^{(1)}$. The distribution of $I_{xy}$ resembles chiral behavior but $I_{xy}$ is homogeneous in each branch (away from the junction region).
In this section, we discuss the Drude model and, using the Landauer-Büttiker formalism (not to be confused with the LBP), derive a coherent and chiral model.
The magnetic field regions in between quantized plateaus deserve special attention. Because, near the plateaus of the QHE, the quantization of the density of states into Landau levels is relevant, we can classify these regions as belonging to the quantum regime. However, between plateaus, the electron transport is diffusive and our coherent and chiral model fails. In fact, the mobility $\mu(B)$ strongly oscillates, such that $\mu B$ can be in the order of 1, as determined from the Shubnikov-de-Haas oscillations of the longitudinal resistance. Can we then apply the Drude model at large magnetic fields for the regions between quantized plateaus and is the magnetotransport between the plateaus chiral or not? In Sec. III we will provide an answer to these questions.
II.1 Drude model for a three-terminal junction
Using $\vec{j}=en_{\text{s}}\vec{\upsilon}_{\text{d}}$ we can rewrite Eq. (1) in terms of the magnetic-field-dependent conductivity tensor as
$\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}=\frac{\sigma_{0}}{1+(\mu B)^{2}}\begin{pmatrix}1&-\mu B\\ \mu B&1\end{pmatrix}\,\begin{pmatrix}E_{x}\\ E_{y}\end{pmatrix}\,,$ (2)
where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$. For a laterally confined 2DES in the steady state, we assume $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$. Following the structure of the conductivity tensor, we devide the current density $j_{x}$ into two components
$j_{x}\equiv j_{xx}+j_{xy}\quad\text{with}$
$j_{xx}=\frac{en_{\text{s}}\mu}{1+(\mu B)^{2}}\,E_{x}\quad\text{and}\quad j_{xy}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{\text{H}}}\,E_{y}\,.$
The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is driven by the field $E_{x}$, which acts along the Hall bar because of its finite resistivity $\sigma_{0}^{-1}$. The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is linked to the Hall voltage and attenuated by the Hall resistance. For $B=0$ the component $j_{xy}=0$, while for $\mu B\to\infty$ the current density becomes independent of $\mu$ as $j_{xx}\to 0$.
[…]
$I=I_{xx}+I_{xy}\quad\text{with}$ (4)
$\frac{I_{xx}}{I}=\frac{1}{1+(\mu B)^{2}}\quad\text{and}\quad\frac{I_{xy}}{I}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,.$
[…]
First, we discuss the current components $I_{xx}$ and $I_{xy}$ for the relatively simple case of a symmetric three-terminal junction as sketched in Fig. 2. We apply a voltage $V_{1}$ to contact 1 while the other two contacts are connected to the electrical ground, $V_{2}=V_{3}=0$. Each contact has an ohmic contact resistance $R_{1,2,3}$ and we apply a finite magnetic field perpendicular to the conducting plane and pointing into it. We can measure the currents $I_{1,2,3}=I_{xx}^{(1,2,3)}+I_{xy}^{(1,2,3)}$. They fulfill Kirchhoffs junction rule $I_{1}+I_{2}+I_{3}=0$ also for the components $I_{xx}^{(1)}+I_{xx}^{(2)}+I_{xx}^{(3)}=0$, $I_{xy}^{(1)}+I_{xy}^{(2)}+I_{xy}^{(3)}=0$, because the respective driving fields are prependicular to each other. The longitudinal component $-I_{xx}^{(1)}$ will be divided between contacts 2 and 3 according to Ohm’s law as indicated in Fig. 2(a): $I_{xx}^{(2,3)}=-I_{xx}^{(1)}\frac{R_{3,2}}{R_{2}+R_{3}}$. The transversal component $-I_{xy}^{(1)}$ follows the charged edge of the Hall bar and is accordingly bent into contact 2 (or into contact 3 for $\vec{B}$ pointing out of the plane). In Fig. 2(b), we indicate the transversal component for the simplified case of a vanishing contact resistance, $R_{2}=0$, resulting in $I_{xy}^{(2)}=-I_{xy}^{(1)}$ and $I_{xy}^{(3)}=0$.
In a realistic device with finite contact resistances, the voltage drop across each contact resistance causes a partial reflection at each contact. Therefore, a smaller fraction of $-I_{xy}^{(1)}$ flows from contact 2 to contact 3 as shown in Fig. 2(c). Because the magnitude of $I_{xy}^{(1)}$ is predetermined by the applied voltage and the Hall resistance, we can express the division of $-I_{xy}^{(1)}$ in terms of two parallel channels with the effective resistances $R_{2}$ versus $R_{\text{H}}+R_{3}$. Hence, we have $I_{xy}^{(2)}=-I_{xy}^{(1)}\frac{R_{\text{H}}+R_{3}}{R_{\text{H}}+R_{2}+R_{3}}$ and $I_{xy}^{(3)}=-I_{xy}^{(1)}\frac{R_{2}}{R_{\text{H}}+R_{2}+R_{3}}$.
The ohmic components $I_{xx}^{(1,2,3)}$ vanish quickly for $\mu B\gg 1$. The transveral currents $I_{xy}^{(1,2,3)}$ vanish at $B=0$, but for $\mu B\gg 1$ the current density distribution merits a closer look. On the one hand, away from a junction, the Drude model predicts a homogeneous current density across a homogeneous sample, cf. Eq. (1), classifying $j_{x}$ to be non-chiral. On the other hand, near a junction, the edge charges caused by the Lorentz force break the homogeneity of the transversal current density $j_{xy}$, while the current $I_{xy}$ is bent into just one of the branches of the junction. Finally, if considering the currents flowing out of each contact in a measurement at $\mu B\gg 1$, practically being a measurement of the transversal components $I_{xy}^{(1,2,3)}$, we expect to find identical result whether the transport is chiral or not. Hence, an experimental indication of non-chiral transport in a multiterminal current measurement can only be expected for $\mu B\lesssim 1$, where the homogeneous $j_{xx}$ is sizable.
II.2 Coherent and chiral model of $n$-terminal device
Figure 3: Simplified circuit diagram of an $n$-terminal device. The inner circle indicates coherent and chiral current flow inside the 2DES. Each contact $i=1,...,n$ has an ohmic resistance $R_{i}$ and a voltage $V_{i}$ applied in respect to the electrical ground. $V_{i}^{\prime}$ are the voltages resulting at the intersections between the ohmic contacts and the coherent inner circle. The magnetic field vector pointing into the plane of the 2DES gives rise to a right handed chirality of a free charge which translates into a left handed chirality of the device, cf. circled arrow.
Next, we derive a general model for the current in a 2DES device prepared in the coherent and chiral limit but equipped with $n$ ohmic contacts, cf. Fig. 3. The model intends to describe current flow through a Hall bar in the region of a quantized plateau. The inner circle in Fig. 3 indicates the coherent 2D conductor, the radial lines symbolize the $n$ terminals $i=1,2,...,n$. External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the coherent part of the sample and each contact. The current $I_{i}$ flows through the $i$th contact with ohmic resistance $R_{i}$ to ground.
To describe the coherent inner part of the sample, we employ the Landauer-Büttiker formalism and write
$I_{i}=\sum_{j=1}^{n}G_{ij}V_{j}^{\prime}=(V_{i}^{\prime}-V_{i})/R_{i}\,,$ (5)
where $G_{ij}$ are the coherent conductance coefficients, which can be formally derived from the scattering matrix of the problem. Assuming chiral transport, current flows only between adjacent contacts. The direction of the chirality is linked to the direction of the magnetic field vector. In our case as sketched in Fig. 3, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}=1$ and $G_{ji}=0$ for $i=j+1$. The unitarity of the scattering matrix implies $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our chiral conductivity tensor takes the form
$G_{ij}=\nu G_{0}\left\{\begin{matrix}1\,;&i=j+1\\[4.30554pt] -1\,;&i=j\hfill\par\LTX@newpage\\[4.30554pt] 0\,;&\text{else}\hfill\par\LTX@newpage\end{matrix}\right.\,,$ (6)
where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels. The coherent limit applies for the quantized plateaus of the Hall resistance, where $\nu$ is an integer. Inserting Eq. (6) into Eq. (5) and solving the equation system yields the currents
$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)$ (7)
where we introduced
$P_{i}=\left\{\begin{matrix}\prod_{j=1}^{i}\left(1+\frac{1}{\nu G_{0}R_{j}}\right)\,;&i=1,2,...,n\\[4.30554pt] 1\,;&i=0\phantom{,2,...,n}\end{matrix}\right.$ (8)
and
$S_{i}=\sum_{j=1}^{i}P_{j-1}\frac{V_{j}}{\nu G_{0}R_{j}}\,.$ (9)
Equation (7) describes the current distribution into the $n$ ohmic contacts of a chiral and coherent sample. For $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$ the currents can be written as
$I_{i>1}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,I_{1}\,,$ (10)
which are independent of $R_{1}$ as expected.
II.3 Drude model of $n$-terminal device
Our model describing coherent currents is based on the unitarity of the scattering matrix. However, regarding the actual macroscopic currents into various contacts, the unitarity condition implies the same result as we find from applying Kirchhoff’s laws. Therefore, if we replace $R_{\text{H}}$ by $(\nu G_{0})^{-1}$, for $n=3$ and $V_{2}=V_{3}=0$, Eq. (7) of the coherent and chiral model predicts the identical currents $I_{2,3}$ as the transversal current component $I_{xy}^{(2,3)}$ of the Drude model discussed in Sec. II.1 above. By extrapolating this observation, we can expand our prediction of the Drude model for the current distribution of an $n$-terminal junction and find
$I_{i}$ $=$ $\frac{1}{1+(\mu B)^{2}}I_{xx}^{(i)}+\frac{(\mu B)^{2}}{1+(\mu B)^{2}}I_{xy}^{(i)}\quad\text{with}$
$I_{xx}^{(i)}$ $=$ $\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (11)
$I_{xy}^{(i)}$ $=$ $\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,.$
The longitudinal component, $I_{xx}^{(i)}$, is determined by Ohm’s law while the transversal component, $I_{xy}^{(i)}$, is identical to Eq. (7), where we use $R_{\text{H}}=(\nu G_{0})^{-1}$.
To correctly describe the regions of the plateaus of a quantized Hall resistance, we use constant and integer $\nu$ and assume $\mu=\infty$, a property of ballistic transport, such that $I_{xx}^{(i)}=0$. Then Eq. (II.3) is identical to Eq. (7).
III experimental results and discussion
Multiterminal current measurements yield the distribution of the current into the different contacts. In this section, we compare the results of multi-terminal measurements and theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$. We will see, that multiterminal measurements can provide answers to the questions about chirality and the local current distribution in magnetotransport.
III.1 Sample and measurement setup
We use a Hall bar as sketched in Fig. 4.
Figure 4: Sketch of a Hall bar with six contacts, characterized by their ohmic resistances, $R_{i}$. We apply a voltage $V_{1}$ to contact 1 and measure currents into contacts 1, 3, and 5. Contacts 2, 4, and 6 can be left floating or connected to electrical ground. Current amplifiers at contacts 3 and 5 result in additional voltages of $\simeq 20\mu$V applied to contacts 3 and 5 (not shown, but taken into account for generating model curves). $\vec{B}$ points into the plane of the 2DES.
It is etched from a GaAs / (Al,Ga)As heterostructure, which contains a two-dimensional electron system (2DES) 130 nm beneath the surface. We performed the experiments at a temperature of $T\simeq 300\,$mK in a He-3 evaporation cryostat. At this temperature and zero magnetic field, the electron density and mobility of the 2DES, determined from Hall measurements, are $n_{\textrm{s}}\simeq 1.2\times 10^{11}\textrm{ cm}^{-2}$ and $\mu\simeq 3.95\times 10^{6}\textrm{cm}^{2}/\textrm{Vs}$ (yielding a mean-free-path of $\lambda_{\text{m}}\simeq 23\,\mu$m), respectively.
Figure 5: Three-terminal measurement with symmetric circuit, cf. inset in (a). A voltage of $V=1\,$mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating. $\vec{B}$ points from the surface into the 2DES plain such that $I_{xy}$ is bent towards contacts 2 and 3. (a) Current ratios $-I_{3}/I_{1}$ and $-I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq. (II.3) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq. (II.3) with $\mu(B)$ as free parameter (dotted red line). (b) enlargements near $\nu=1$ showing experimental data and the Drude model prediction for $\mu=\mu_{0}$.
Figure 6: Mobility $\mu(B)$ as determined from SdH-oscillations of the longitudinal resistance between contacts 2 and 3 (black dots) in comparison with $\mu(B)$ found by fitting Eq. (II.3) to the three-terminal measurements shown in Fig. 5 (red squares).
Figure 7: Four-terminal measurements for three different contact configurations, cf. insets, further configurations and labeling as in Fig. 5. (a) symmetric four-terminal circuit. (b) and (c) variations of asymmetric four-terminal circuits.
For multi-terminal current measurements, we apply the voltage of $V_{1}=\pm 1\,$mV to the source contact 1 (using a Keithley 2450 source meter). We measure the current $I_{1}(B)$ flowing from the Hall bar into contact 1, using the source meter, and the currents $I_{3}(B)$ and $I_{5}(B)$ flowing into contacts 3 or 5, respectively, using Basel Precision Instr. current amplifiers. Contacts 3 and 5 are connected to the electrical ground, while the remaining contacts 2, 4, and 6 may be either electrically floating or connected to ground depending on the experiment, cf. Fig. 4. Here, we apply $B$ perpendicular to the 2DES, such that it points from the sample surface towards the 2DES, hence, electrons contributing to $I_{xy}$ move from contact to contact along the counterclockwise pathway $1\to 2\to 3\to 4\to 5\to 6\to 1$. Due to the direction of the magnetic field and a voltage applied to contact 1, the largest portion of $I_{xy}$ will always flow between contact 1 and contact 3, if contact 2 is floating, or to contact 2, if it is grounded. Whether this current is positive or negative depends on the sign of $V_{1}$.
We performed a series of measurements at $B=0$ to determine the contact resistances ($R_{i}\sim 2.5\,\text{k}\Omega$) and offset voltages ($V_{3,5}\sim 20\,\mu$V) of the two current amplifiers. In addition, we used the comparison between measured data and model curves for a fine tuning of these circuit parameters, which remain constant between measurements. They are listed in Table 1.
contact $i$ resistance $R_{i}$ ($\Omega$) voltage $V_{i}$ (mV) geometry $a_{i}$
1 2420 1.000 1
2 2420 0 0.5
3 2400 -0.021 1
4 2350 0 1
5 2420 -0.018 1
6 2400 0 0.5
Table 1: Parameters of the circuit sketched in Fig. 4. Contact resistances $R_{i}$ include ohmic contacts, cables and $RC$-filters. Voltage $V_{1}$ is applied to contact 1, $V_{2}$ and $V_{3}$ are offset voltages of the current amplifiers. To account for the contribution of the Hall bar itself to the resistance, we defined a geometry parameter and use $R_{i}\to R_{i}+a_{i}/[n_{\text{s}}e\mu(B)]$ in the model calculations of the ohmic contributions. The correction is negligible for $B=0$.
III.2 Three-terminal measurements
First, we discuss a three-terminal measurement in a mirror symmetric geometry, for which we keep the contacts 2, 4 and 6 electrically floating; current flows from contact 1 into contacts 3 and 5. In Fig. 5(a), we present as solid (blue) lines the measured current ratios $-I_{3}(B)/I_{1}(B)$ and $-I_{5}(B)/I_{1}(B)$. As expected, we find that $|I_{3}|$ increases with $B$ while $|I_{5}|$ decreases accordingly, such that $I_{3}(B)+I_{5}(B)=-I_{1}(B)$. The dashed black line indicates the prediction of the Drude model according to Eq. (II.3) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$. For this large mobility, the Ohmic term (first term) of Eq. (II.3) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$). In this limit, with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$, Eq. (II.3) becomes identical to the coherent chiral model expressed in Eq. (7).
The Drude model and the measured data agree in the classical limit for $B\lesssim 0.5$ T. As we increase $B$ further, the Drude model with a constant $\mu(B)=\mu_{0}$ still predicts the correct trend, but we observe growing deviations for filling factors away from integer values [top axis of Fig. 5(a)]. The enlargement around the first plateau of the QHE in Fig. 5(b) clarifies, that model and experiment match exactly only at specific values of $B$ including integer bulk filling factors (here $\nu=1$). An exact match between measurements and model can be achieved by using $\mu(B)$ as a free fit parameter in Eq. (II.3). For the dotted red line in Fig. 5(a), we used the magnetic field dependent values of $\mu(B)$ as shown by red squares in Fig. 6. For comparison, we have also determined $\mu(B)$ from the Shubnikov-de-Hass oscillations of a standard longitudinal resistance measurement, see the black dots in Fig. 6.  [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to $T>4\,$K between the two measurements. The agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq. (II.3) to our three-terminal current measurements supports the validity of our model.
The two regions of divergent $\mu(B)$ in Fig. 6 correspond to the plateau regions at $\nu=2$ and $\nu=1$. Here, $I_{xx}=0$ and the coherent and chiral model expressed in Eq. (7) is valid. Away from the plateau regions, we find strongly reduced mobility values with $201\lesssim\mu<$ 20, which indicate diffusive transport and a finite ohmic contribution $I_{xx}$ to the current.
The overall agreement between the prediction of Eq. (II.3) and our three-terminal measurements is convincing. On the one hand, it supports the assumption of coherent and chiral transport for the quantized plateaus of the Hall resistance. On the other hand, it points to non-chiral transport described by the Drude model not only in the classical regime of the Hall effect but also for large $B$ in between the plateaus.
III.3 Four-terminal measurements
In Fig. 7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements. Again, we measure the currents into contacts $i=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig. 7(a), contact 6 in (b), and contact 2 in (c). If we assume $\mu(B)=\mu_{0}$, Eq. (II.3) of the Drude model reproduces the overall dependences $I_{i}(B)$ with a similar accuracy as it did for the three-terminal measurement (dashed black lines). The dotted red lines correspond to the predictions of Eq. (II.3) using the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf. Fig. 6. Again, for $B\lesssim 0.5\,$T, theory and experiment agree well. For $B>0.5\,$T, the agreement is still perfect for the plateau regions but becomes worse away from the plateaus of quantized Hall resistance. experimental results is even worse (for $B>0.5\,$T).
This observation allows two immediate conclusions: (1) for the plateau regions, the chiral transport already found in the three-terminal measurement is confirmed; (2) away from the plateaus, the non-chiral transport is also confirmed. However, for more than three terminals, the model prediction of Eq. (II.3) fails for $B>0.5\,$T.
In Fig. 8,
Figure 8: Comparison of the respective largest measured current contributions flowing into a grounded contact for three different configurations of four-terminal measurements. The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled such that all three theory curves according to Eq. (II.3) with $\mu=\mu_{0}$ are identical following the edge between shaded and white backgrounds.
we plot for each of the three four-terminal measurements shown in Fig. 7 the branch corresponding to the largest current. We slightly scaled the data to compensate for small variations in the contact resistances. After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) and on the plateaus of the QHE. The edge between shaded and white backgrounds corresponds to the prediction of the Drude model for constant $\mu=\mu_{0}$ [dashed black lines in Fig. 7(a)]. Compared to this model prediction, the current is reduced between the plateaus for all three geometries. The reduction is particularly strong for the set-up with contact 6 connected to ground. In the case of chiral transport, contact 6 is the last one to be reached by the current, while most current should flow into contact 3. The reduction of the current into contact 3 goes along with an enhancement of the current into contact 6, cf. Fig. 7(b). This is a clear signature of non-chiral current but, the current reduction into contact 3 (and the current enhancement into contact 6) exceeds even the prediction of our model expressed by Eq. (II.3). Likewise, for the set-up with contact 2 grounded, the reduction of the current into contact 2 is smaller than the reduction predicted by Eq. (II.3). So, a smaller amount of the non-chiral current reaches contact 5 as it is much further away from contact 1 than contact 2.
The details of the measured current distribution clearly depend on the contact configuration. This behavior demonstrates the limits of our model. For the derivation of Eq. (II.3), we assumed an infinitely long Hall bar without additional grounded contacts. Hence, for identical contact resistances, its prediction is independent of the variations between specific configurations of the four-terminal set-up (the resistance of the Hall bar itself at $B=0$ is much smaller than the contact resistances). To overcome the failure of our simplified model, it would be necessary to account for the more complicate steady-state electrostatics of a multi-terminal device, in particular the inhomogeneity of the electric field near a junction. Suitable numerical calculations could clarify this point. For the moment, we conjecture that such a corrected model would support our previous result based on the three-terminal measurement, namely that the Drude model correctly describes the transition regions between plateaus.
IV Screening theory and chiral transport
Figure 9: Coherent and chiral models for the first plateau of the QHE: Sketch of the distribution of the current $I_{xy}$ near $\nu=1$ for $V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{2}=R_{3}$ with $\vec{B}$ pointing into the plane of the 2DES. Color intensities indicate the distribution of the absolute value of the current density into the three branches. Black bent arrows show the direction and relative magnitudes of electron flow related with the charged edges indicated by blue minus signs for $V_{1}<0$. For $V_{1}>0$, current densities remain the same while the edge charges become positive and black arrows indicate the current flow. (a) Prediction of the Landauer Büttiker picture, assuming that current flows in one-dimensional chiral edge channels only. (b) Prediction of the screening theory for the case of edge ICSs (current flows inside ICSs only). (c) Prediction of the screening theory for the case of a bulk ICS. All three cases predict chiral current flow within the 2DES and result in identical currents $I_{1}$, $I_{2}$ and $I_{3}$ into the contacts. In the stem towards contact 1 we sketch the electrochemical potentials of the first Landau level (LL1, solid line) and the related Hall potential ($U_{\text{H}}$, dotted red line).
In this section, we restrict the discussion to the plateau regions of the quantized Hall resistance, for which our multiterminal experiments indicate chiral and coherent transport. The unique characteristics of chiral transport of a 2D conductor is the opposite directions of current flow along the opposite edges of the conductor. The QHE reveals the additional feature of fundamental resistance values being fractions of $h/e^{2}$. These values can be measured with an extremely high accuracy and stability with deviations of less than $3\times 10^{-9}$ having been demonstrated between high resolution measurements of various samples in different laboratories [16]. In order to explain the measured highly accurate resistance values in fractions of $h/e^{2}$, scattering has to be (almost) completely suppressed.
The LBP [1, 2] is a single-particle model based on the assumption that the Landau levels form perfect 1D edge channels, which support chiral transport, while the bulk of the 2DES is assumed to be electrically insulating because of disorder-induced localization. The edge channels are compressible, meaning that there exist occupied and unoccupied states at the chemical potential. The mere assumption of scattering free transport via compressible edge channels as an explanation of the extremely accurate measurements of the fundamental resistance value $h/e^{2}$ is unsatisfactory.
The screening theory offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$, which is based on first principle calculations [7, 8, 9]. Already the earliest purely classical considerations, taking into account the Coulomb interaction, led to the realization of a segmentation of the 2DES confined in a Hall bar in perfectly screened compressible regions and completely unscreened ICSs [3, 4, 5, 6]. Inside the ICSs, scattering is forbidden because the excitation of a carrier costs the energy $\hbar\omega_{\text{c}}$ of the Landau gap. While numerous experiments already support the screening theory [17, 10, 11, 12, 18, 19, 20, 21, 22, 23, 13], here our focus is on the question whether the screening theory is consistent with chiral transport.
In Fig. 9, we sketch the situation for the first plateau ($\nu=1$, $I_{xx}=0$) of the QHE for the case of a three-terminal junction. As a start, in Fig. 9(a), we show the situation assumed in the LBP. All current flows inside the compressible edge channels (solid blue lines). The stem towards contact 1 contains a sketch of the first Landau level (LL1, solid line), which strictly follows the confinement potential for the single particle LBP. The edge channels emerge, where LL1 intersects with the chemical potential, which differs by $V_{\text{H}}$ between the respective edges. The Hall potential $U_{\text{H}}=eV_{\text{H}}$ (dotted line) drops linearly between the edge channels.
To illustrate the screening theory [9], we consider the two limiting cases realized near the ends of the first plateau of the QHE. For simplicity, we neglect disorder effects. Near the lower magnetic field end with $\nu>1$, the screening theory predicts two narrow ICSs roughly resembling Landauer-Büttiker edge channels, cf. Fig. 9(b). Near the high field end of the plateau, where $\nu\leq 1$, the prediction is a single bulk ICS, while the 2DES near the edges remains compressible, cf. Fig. 9(c). Even for the first case with two edge ICSs, in contrast to the LBP, the current flows entirely inside the ICSs, where scattering is suppressed due to the Landau gap of the density of states. In the compressible regions of the 2DES, electrons can scatter and the local electric field is perfectly screened. As a result, the local electric field entirely drops inside the ICSs. The gradient of the Landau levels, i.e., the effective local electric field drives the current according to $\vec{j}=-en_{\text{s}}\,\vec{E}\times\vec{B}/B^{2}$. Therefore, no current flows in the perfectly screened compressible regions. The gradient of the potential energy of the Landau levels equals the gradient of the local electric field. As can be seen in the sketches of the first Landau level in Figures 9b and 9c, the gradient of the Landau levels has opposite directions inside the ICS(s) along the oposite edges of the sample. This is even the case if only a single ICS covers most of the 2DES, cf. Fig. 9(c). Here, $\vec{E}=0$ and $\vec{j}=0$ in the center of the ICS and the current flows in opposite directions on both sides of the ICS. Consequently, even in case of a single bulk ICS, the current flow is still chiral and the local current density is larger near the sample edges, where the confinement potential drops steeply.
V Summary and Outlook
In summary, we have performed multi-terminal current measurements in the regime of the QHE and compared our results with predictions of a coherent and chiral model versus the Drude model. The Drude model perfectly describes our results in the classical regime of small magnetic fields, the coherent and chiral model fits for the regions corresponding to quantized plateaus of the Hall resistance. In the diffusive regions for large magnetic fields, i.e., between quantized plateaus our three-terminal experiment supports the predictions of the Drude model, which includes non-chiral transport. The four-terminal measurements demonstrate the limits of our model but clearly demonstrate non-chiral current contributions away from the plateaus of quantized Hall resistance. This work contributes to a deeper understanding of the microscopic dynamics and the distribution of current density inside realistic mesoscopic devices in the regime of the QHE.
Acknowledgement
The authors thank Piet Brouwer for help with the coherent model and Afif Siddiki for fruitful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 218453298.
Contributions of the authors
V. Y. U. supplied the wafer. S. S. and M. K. performed the measurements. S. S. and S. L. analyzed the data. S. L. derived the models. S. S. and S. L. wrote the article.

### v2

@@
-Current distribution and chirality in the regime of the quantized Hall effect
+Chiral nature of current flow in the regime of the quantized Hall effect
 Abstract
-The integer quantum Hall effect (QHE) is one of the most basic and, at the same time, astonishing phenomena in condensed matter physics.
-Despite its importance for metrology and fundamental physics, our microscopic understanding of the QHE is not yet complete.
-For decades, there has been a discussion about the current distribution inside a Hall bar for the case of a quantized Hall resistance.
-With the advent of quantum technology, the questions, if the current flows in compressible Landauer-Büttiker edge channels or in incompressible strips with varying geometry and whether the current flow is chiral or not, gain importance, because the quantum phase of each charge carrier depends on its exact pathway.
-In the present work, we compare multi-terminal current measurements in the regime of the QHE with model calculations in order to unravel the current density distribution in a Hall-bar device.
-We find, that the conjecture of chirality is justified for the plateaus of quantized Hall resistance, while in between plateaus the current density looses its chirality.
-The latter is described by the Drude model, which predicts a homogeneous current density for a sample with homogeneous mobility and carrier density.
+The integer quantum Hall effect (QHE) belongs to the most fundamental phenomena of solid state physics and has an important application as resistance standard.
+It serves as a basis to understand the fractional, anomalous or spin QHEs, candidates for applications in quantum technology due to their topological properties.
+For optimizing all these applications it is essential to understand the current distribution inside the Hall bar, a question disputed for decades.
+We perform multiterminal current measurements on a Hall bar and compare the results with limiting models.
+We show, based on these simple experiments, that the current flow is chiral for the plateaus of quantized Hall resistance.
+Everywhere else, the current is described by the Drude model.
 I Introduction
@@
 This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$.
-The axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ in the classical equation of motion $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$ of an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$.
-As an introduction, we discuss the dynamics of charge carriers considering a variety of boundary conditions.
-The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a constant drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf.
+In this article, we explore the implication of the axial symmetry of $\vec{B}$ for the current flow in the regime of the quantized Hall effect.
+In the classical equation of motion, $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$, the axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ acting on an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$.
+The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf.
 Fig.
 1(a).
-The chirality of this trajectory is left-handed for the negatively charged electron (but would be right handed for a positive charge).
-Note that we define the chirality of a trajectory in analogy to the handedness of a screw thread.
+If we define the chirality of a trajectory in analogy to the handedness of a screw thread, the above trajectory of a ballistic electron is left-handed (and would be right handed for a positive charge).
 Figure 1: Motion (expressed by drift velocity) of an electron confined to the $xy$-plane with in-plane electric and perpendicular magnetic fields, $\vec{E}$ and $\vec{B}$ (red arrows), respectively.
 (a) Ballistic limit (free electron).
-The spiral is the classical trajectory consisting of a constant drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity.
+The spiral is the classical trajectory consisting of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity.
 (b) Diffusive limit.
@@
 Its key parameter is the electron mobility $\mu$.
-For $\mu B<1$ (corresponding to a mean-free path $l_{\text{m}}<R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering.
-Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$ and applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$, where $\tau=m\mu/e$ is the momentum scattering time.
-In this diffusive limit and for $\vec{B}\perp\vec{E}$, the solution of the equation of motion introduced above can be written as
-$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,.$ (1)
-[For $\vec{E}\not\perp\vec{B}$ there would be a third summand, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$, inside the parantheses.] Clearly, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous.
-However for finite $B$, $\vec{j}$ is not parallel to $\vec{E}$, but bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$, cf.
+For $\mu B\lesssim 1$ (corresponding to a mean-free path $l_{\text{m}}\lesssim R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering.
+Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$.
+Applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$ with the momentum scattering time $\tau=m\mu/e$, the solution of the equation of motion introduced above in the diffusive limit becomes
+$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,\text{ for }\vec{B}\perp\vec{E}\,.$ (1)
+[For $\vec{B}\not\perp\vec{E}$ there would be a third summand inside the parantheses, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$.] For finite $B$, $\vec{j}$ is bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$ in respect to $\vec{E}$, cf.
 Fig.
 1(b).
-The ballistic limit discussed above is clearly beyond the scope of the Drude model.
-Indeed, for $\mu B\to\infty$, the friction forces in Eq.
-(1) become negligible compared to the Lorentz force and the drift velocity becomes independent of $\mu$.
-Nevertheless for $\mu B\to\infty$, the drift velocity takes the value $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, which is identical to that of a ballistic electron.
-This observation encourages us to compare measurements with the Drude model even for $\mu B\gg 1$.
-In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, and include boundaries of the 2DES and electron scattering.
-Furthermore, we consider an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular, homogeneous magnetic field $\vec{B}=(0,0,B)$, corresponding to one of the most common set-ups for magnetotransport measurements.
-For simplicity, we restrict our general discussion to a homogeneous electron density and mobilty inside the 2DES (away from the edges).
+Importantly, as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous, too.
+While the Drude model is designed to describe the limit of $\mu B\lesssim 1$, its prediction for the drift velocity for $\mu B\to\infty$ where the friction forces in Eq.
+(1) become negligible, namely $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, is identical to the ballistic solution.
+Since for $\mu B\to\infty$ the rapid cyclotron motion averges out over reasonable length scales, both models predict identical current densities, too, for $\mu B\to\infty$.
+In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, with an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular homogeneous magnetic field $\vec{B}=(0,0,B)$.
+This is a common scenario realized in many mesoscopic magnetotransport measurements.
 To discuss the classical limit of the Hall effect, we include two sample edges at $y=0$ and $y=b$, such that the electrons are confined in the $y$-direction but move freely in the $x$-direction, cf.
@@
 In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons.
-Applying this condition to Eq.
-(1) we find $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$.
-The transversal field $E_{y}=\vec{\nabla}V_{\text{H}}$ is maintained by charge accumulation at the sample edges, which gives rise to the Hall voltage $V_{\text{H}}$.
-In the steady state, $-eE_{y}=-F_{\text{L}}$, i.e., the Lorentz force acting on a free electron is exactly canceled, and the electrons move as if no magnetic field was applied but only the longitudinal field $E_{x}$.
-For our homogeneous sample, the current density $j_{x}$ is a constant and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
-In a confined sample, the axial character of the magnetic field manifests itself in the edge charges or the related Hall voltage (which break the time reversal symmetry).
-The latter gives rise to an increase of the two-terminal resistance of the 2DES according to $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}$ with the resistance $R_{0}$ of the 2DES at $B=0$.
-As we increase $B$, the quantization of the density of states into Landau-levels results in the quantized Hall effect (QHE).
-Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance.
-The latter indicates scattering-free transport within the Hall bar.
-To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of perfect one-dimensional (1D) edge channels, in which scattering is suppressed.
-The conductance of a spin-resolved 1D channel is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$.
-The filling factor $\nu$ describes the number of perfect 1D channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk.
-The perfect 1D-channels are assumed to follow the edges of the Hall bar, because they form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2].
-To obtain a quantized Hall resistance, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization.
+Applied to Eq.
+(1) $j_{y}=0$ yields $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$.
+Physically, the transversal field $E_{y}$ is sourced by charge accumulated at the sample edges.
+With $E_{y}=\vec{\nabla}V_{\text{H}}$ it also gives rise to the Hall voltage $V_{\text{H}}$, that can be measured across the Hall bar.
+Because in steady state, the Lorentz force, acting on the carriers, is exactly canceled by $-eE_{y}$ (the forces associated with $B$ and $E_{y}$ cancel each other), inside the Hall bar the electrons move as if only the longitudinal electric field $E_{x}$ was applied.
+Given a homogeneous sample (homogeneous distributions of $n_{\text{s}}$ and $\mu$ away from the edges of the Hall bar), the current density $j_{x}$ is uniform and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
+While the Drude model predicts a homogeneous current density for the classical Hall effect, the axial character of the magnetic field and the breaking of the time-reversal symmetry are manifest in the edge charges and the related Hall voltage.
+In addition, the Hall resistance is non-local as it does not cause any dissipation inside the Hall bar: At $B=0$ the two terminal current measurement of a Hall bar is the sum of the ohmic sample resistance and the resistance of the leads (including the ohmic contacts), $R_{\text{2T}}(0)=R_{0}+R_{\textrm{leads}}$.
+At finite $B$, the non-local Hall resistance enhances the resistance of the Hall bar, such that $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}+R_{\textrm{leads}}$.
+As we increase $B$, at low enough temperatures the quantization of the density of states into discrete Landau-levels results in the quantum Hall effect (QHE).
+Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance ($R_{0}=0$).
+The latter indicates the absence of momentum scattering within the Hall bar for a quantized Hall resistance, in agreement with the classical Drude prediction for $\mu B\to\infty$.
+To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of chiral, one-dimensional (1D) edge channels, such that backward scattering is suppressed [1, 2].
+The conductance of a spin-resolved 1D channel without momentum scattering is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$.
+The filling factor $\nu$ describes the number of edge channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk.
+The LBP assumes that the edge channels form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2].
+To obtain finite ranges of quantized Hall resistances, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization.
 The direction of the current flow in the edge channels is predetermined by the gradient of the confinement potential.
-This yields current flow in opposite directions at opposite edges.
-In that sense, in the LBP the QHE is associated with chiral current flow.
+This yields chiral current flow, more precisely, current flow in opposite directions along opposite edges.
 The LBP provides an intuitive explanation for the quantized Hall resistance and the vanishing longitudinal resistance of the QHE.
@@
 In particular, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [3, 4].
-However, this would cause highly charged stripes between the edge channels corresponding to an unrealistically large Coulomb energy.
-The screening theory [3, 4, 5, 6, 7, 8, 9] overcomes these limitations by taking into account the direct Coulomb interaction between the electrons.
+Hence, neglecting the Coulomb interaction between carriers in the LBP causes the formation of highly charged stripes between the edge channels.
+This local polarization corresponds to an unrealistically large Coulomb energy.
+Another weak point of the LBP is the coexistence of occupied and unoccupied states inside the current carrying edge channels at the chemical potential.
+Even, if we assume chiral edge channels, an almost complete suppression of carrier scattering, as indicated by the accuracy of the measurements of the von Klitzing constant, is unlikely under any realistic conditions.
+The screening theory [3, 4, 5, 6, 7, 8, 9] offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$.
+By taking into account the direct Coulomb interaction between electrons, it overcomes the limitations of the LBP.
 The Coulomb interaction causes a rearrangement of the free carriers such that the highly charged stripes are avoided.
-Instead, it leads to a segmentation of the Hall bar in compressible regions and incompressible strips (ICSs).
-Inside the ICSs, carriers cannot scatter and screening of the electric field is absent, because the chemical potential is in a local energy gap between filled and empty Landau levels; ICSs are accompanied by adjacent compressible regions, in which a Landau-level is aligned with the chemical potential, such that carriers can scatter.
-Inside the compressible regions, the electric field is then perfectly screened.
-The current flows where the electric field drops, i.e., inside the ICSs.
-In contrast to the 1D Landauer-Büttiker edge channels, the ICSs have a finite width exceeding the magnetic length [7, 8, 9].
-Width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$.
-In particular, towards the end of each plateau with increasing magnetic field, the screening theory predicts a single extended incompressible region in the center of the Hall bar [7, 8, 9].
-This implies that current can flow in the bulk of the Hall bar and not just in edge channels.
-The predicted bulk ICS near the high magnetic field end of the plateaus was experimentally corroborated in scanning gate spectroscopy experiments [10, 11, 12] and recently directly confirmed in transport experiments [13].
-As we will discuss in Sec.
-IV below, the screening theory is consistent with chiral transport at the plateaus of the QHE, even if the current flows inside a single bulk ICS.
-In the present article, we investigate the influence of a magnetic field on the current density distribution of a multiterminal junction.
+Early purely classical formulations considered this screening based on the Thomas-Fermi approximation [3, 4].
+They demonstrated that the Landau-quantization leads to a segmentation of the 2DES of a Hall bar in compressible regions and incompressible strips (ICSs).
+The term ICS denotes areas of the 2DES, where the chemical potential lies within a local energy gap between completely filled and empty Landau levels.
+For this reason, at low enough temperature, inside the ICSs carriers cannot scatter and screening of the electric field is absent.
+ICSs are accompanied by adjacent compressible regions, in which a partly filled Landau-level is aligned with the chemical potential.
+Hence, inside compressible regions, carriers can scatter.
+Provided the vicinity of unscreened ICSs, the screening of the local electric field becomes perfect inside the compressible regions.
+The early calculations already show, that both, width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$.
+Later, for given boundary conditions the number and geometry of the ICSs were predicted more accurately using a self-consistent numerical calculation of Poisson’s equation within the Thomas-Fermi approximation and accounting for the quantum mechanical wavefunctions of the electrons in a mean field approach [7, 8, 9] (later confirmed in a full self-consistent approach of the quantum- electrostatic problem [10]).
+The calculations predict, that towards the higher magnetic field end of each plateau the ICSs merge into a single extended incompressible region centered in the Hall bar [7, 8, 9].
+These predictions clearly confirmed the electrostatic landscape measured in scanning probe spectroscopy experiments as a function of the magnetic field [11, 12, 13].
+According to the screening theory the electric field vanishes inside the perfectly screened compressible regions while it is exclusively contained inside the ICSs.
+For suppressed scattering the drift velocity of the electrons is $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$.
+Hence, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, vanishes in the compressible regions, while all current flows inside the ICSs [4, 14, 9, 15].
+In case of a bulk ICS, current should then flow in the center of the Hall bar.
+Recently, we confirmed the predicted bulk current directly by means of an additional tiny ohmic contact in the center of a Hall bar [16].
+Both, the electrostatics and the existence of bulk currents, predicted by the screening theory [9] and confirmed in measurements [11, 12, 13, 16] are in direct contrast to the assumptions of the LBP.
+Therefore, it is important to find out, whether another assumption of the LBP, which is often used, remains correct, namely whether the current flow is chiral for the quantized plateaus of the Hall resistance.
+Even within the screening theory this is still a controversal question [3, 4, 9, 10, 15].
+In the present article, we explore the question of chirality experimentally by employing multiterminal current measurements.
+Thereby, we probe the division of the current into various ohmic contacts of the Hall bar.
+Clearly, we do not directly probe the local current density distribution.
+Nevertheless, it is possible to distinguish between chiral versus non-chiral transport as the current division into multiple contacts directly depends on it.
+To quantify this dependence, we use the properties of the macroscopic scattering matrix, which yields the division of the current into various contacts as follows: In the case of chiral current flow, only unidirectional transmission coefficients connecting adjacent contacts are non-zero.
+In comparison, non-chiral current flow requires additional non-zero transmission coefficients, which would ultimately yield a different divison of the current into the various contacts.
+Here, we show that the measured current division can be described for arbitrary magnetic fields by either one of two limiting models, introduced in Sec.
+II below.
 In Sec.
-II we illuminate the non-chiral character of the Drude model and derive a general coherent and chiral model.
-In Sec.
-III we compare our multi-terminal current measurements with the predictions of the models.
-For relatively small magnetic fields, we find that the classical limit of the Hall effect is described by the Drude model.
-For higher magnetic fields, we distinguish between the plateau regions of a quantized Hall resistance and the transition regions between plateaus.
-For the plateaus, coherent and chiral transport was already supported by Mach-Zehnder interferometer experiments measuring interference between two distinct current paths [14].
-Our own results confirm this.
-For the magnetic field regions in between the plateaus our measurements are consistent with the predictions of the classical Drude model, which predicts non-chiral transport.
-In Sec.
-IV we explain, why the screening theory predicts chiral transport.
-II model
-Figure 2: Drude model: Distribution of current $I=I_{xx}+I_{xy}$ in a three-terminal device with homogeneous electron density and mobility in the diffusive and classical limit for $V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{2}=R_{3}$ and $\vec{B}$ pointing into the plane of the 2DES.
-Color intensities indicate the distribution of the absolute value of the current density into the three branches.
-Within each branch, the current density is homogeneous.
-For $V_{1}<0$, white arrows show the directions of the conventional currents and black bent arrows the direction and relative magnitudes of electron flow (for $V_{1}>0$ vice versa).
-(a) $I_{xx}$ displaying ohmic behavior, the same as for $B=0$: $I_{xx}^{(2)}=I_{xx}^{(3)}=-0.5I_{xx}^{(1)}$.
-(b) $I_{xy}$ for $R_{2}=R_{3}=0$ displaying non-ohmic behavior: $I_{xy}$ is bent by the Lorentz force: $I_{xy}^{(2)}=-I_{xy}^{(1)}$ and $I_{xy}^{(3)}=0$.
-(c) $I_{xy}$ for $R_{2}=R_{3}>0$: $I_{xy}$ is still bent by the Lorentz force but partly reflected at the resistive contacts: $I_{xy}^{(2)}>I_{xy}^{(3)}$ with $I_{xx}^{(2)}+I_{xx}^{(3)}=-I_{xx}^{(1)}$.
-The distribution of $I_{xy}$ resembles chiral behavior but $I_{xy}$ is homogeneous in each branch (away from the junction region).
-In this section, we discuss the Drude model and, using the Landauer-Büttiker formalism (not to be confused with the LBP), derive a coherent and chiral model.
-The magnetic field regions in between quantized plateaus deserve special attention.
-Because, near the plateaus of the QHE, the quantization of the density of states into Landau levels is relevant, we can classify these regions as belonging to the quantum regime.
-However, between plateaus, the electron transport is diffusive and our coherent and chiral model fails.
-In fact, the mobility $\mu(B)$ strongly oscillates, such that $\mu B$ can be in the order of 1, as determined from the Shubnikov-de-Haas oscillations of the longitudinal resistance.
-Can we then apply the Drude model at large magnetic fields for the regions between quantized plateaus and is the magnetotransport between the plateaus chiral or not?
-In Sec.
-III we will provide an answer to these questions.
-II.1 Drude model for a three-terminal junction
-Using $\vec{j}=en_{\text{s}}\vec{\upsilon}_{\text{d}}$ we can rewrite Eq. (1) in terms of the magnetic-field-dependent conductivity tensor as
-$\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}=\frac{\sigma_{0}}{1+(\mu B)^{2}}\begin{pmatrix}1&-\mu B\\ \mu B&1\end{pmatrix}\,\begin{pmatrix}E_{x}\\ E_{y}\end{pmatrix}\,,$ (2)
-where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$.
-For a laterally confined 2DES in the steady state, we assume $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
-Following the structure of the conductivity tensor, we devide the current density $j_{x}$ into two components
-$j_{x}\equiv j_{xx}+j_{xy}\quad\text{with}$
-$j_{xx}=\frac{en_{\text{s}}\mu}{1+(\mu B)^{2}}\,E_{x}\quad\text{and}\quad j_{xy}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{\text{H}}}\,E_{y}\,.$
-The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is driven by the field $E_{x}$, which acts along the Hall bar because of its finite resistivity $\sigma_{0}^{-1}$.
-The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is linked to the Hall voltage and attenuated by the Hall resistance.
-For $B=0$ the component $j_{xy}=0$, while for $\mu B\to\infty$ the current density becomes independent of $\mu$ as $j_{xx}\to 0$.
-[…]
-$I=I_{xx}+I_{xy}\quad\text{with}$ (4)
-$\frac{I_{xx}}{I}=\frac{1}{1+(\mu B)^{2}}\quad\text{and}\quad\frac{I_{xy}}{I}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,.$
-[…]
-First, we discuss the current components $I_{xx}$ and $I_{xy}$ for the relatively simple case of a symmetric three-terminal junction as sketched in Fig.
-2.
-We apply a voltage $V_{1}$ to contact 1 while the other two contacts are connected to the electrical ground, $V_{2}=V_{3}=0$.
-Each contact has an ohmic contact resistance $R_{1,2,3}$ and we apply a finite magnetic field perpendicular to the conducting plane and pointing into it.
-We can measure the currents $I_{1,2,3}=I_{xx}^{(1,2,3)}+I_{xy}^{(1,2,3)}$.
-They fulfill Kirchhoffs junction rule $I_{1}+I_{2}+I_{3}=0$ also for the components $I_{xx}^{(1)}+I_{xx}^{(2)}+I_{xx}^{(3)}=0$, $I_{xy}^{(1)}+I_{xy}^{(2)}+I_{xy}^{(3)}=0$, because the respective driving fields are prependicular to each other.
-The longitudinal component $-I_{xx}^{(1)}$ will be divided between contacts 2 and 3 according to Ohm’s law as indicated in Fig.
-2(a): $I_{xx}^{(2,3)}=-I_{xx}^{(1)}\frac{R_{3,2}}{R_{2}+R_{3}}$.
-The transversal component $-I_{xy}^{(1)}$ follows the charged edge of the Hall bar and is accordingly bent into contact 2 (or into contact 3 for $\vec{B}$ pointing out of the plane).
-In Fig.
-2(b), we indicate the transversal component for the simplified case of a vanishing contact resistance, $R_{2}=0$, resulting in $I_{xy}^{(2)}=-I_{xy}^{(1)}$ and $I_{xy}^{(3)}=0$.
-In a realistic device with finite contact resistances, the voltage drop across each contact resistance causes a partial reflection at each contact.
-Therefore, a smaller fraction of $-I_{xy}^{(1)}$ flows from contact 2 to contact 3 as shown in Fig.
-2(c).
-Because the magnitude of $I_{xy}^{(1)}$ is predetermined by the applied voltage and the Hall resistance, we can express the division of $-I_{xy}^{(1)}$ in terms of two parallel channels with the effective resistances $R_{2}$ versus $R_{\text{H}}+R_{3}$.
-Hence, we have $I_{xy}^{(2)}=-I_{xy}^{(1)}\frac{R_{\text{H}}+R_{3}}{R_{\text{H}}+R_{2}+R_{3}}$ and $I_{xy}^{(3)}=-I_{xy}^{(1)}\frac{R_{2}}{R_{\text{H}}+R_{2}+R_{3}}$.
-The ohmic components $I_{xx}^{(1,2,3)}$ vanish quickly for $\mu B\gg 1$.
-The transveral currents $I_{xy}^{(1,2,3)}$ vanish at $B=0$, but for $\mu B\gg 1$ the current density distribution merits a closer look.
-On the one hand, away from a junction, the Drude model predicts a homogeneous current density across a homogeneous sample, cf.
-Eq.
-(1), classifying $j_{x}$ to be non-chiral.
-On the other hand, near a junction, the edge charges caused by the Lorentz force break the homogeneity of the transversal current density $j_{xy}$, while the current $I_{xy}$ is bent into just one of the branches of the junction.
-Finally, if considering the currents flowing out of each contact in a measurement at $\mu B\gg 1$, practically being a measurement of the transversal components $I_{xy}^{(1,2,3)}$, we expect to find identical result whether the transport is chiral or not.
-Hence, an experimental indication of non-chiral transport in a multiterminal current measurement can only be expected for $\mu B\lesssim 1$, where the homogeneous $j_{xx}$ is sizable.
-II.2 Coherent and chiral model of $n$-terminal device
+III, we then present our three- and four-terminal current experiments, compare them with the predictions of the two limiting models and discuss our results.
+II models
+Our first model generically assumes coherent and chiral current flow.
+The condition of coherence is technical.
+In practice, it is sufficient to assume the absence of scattering.
+The second model applies the classical Drude model to a multiterminal current measurement.
+It describes diffusive and non-chiral current.
+Our experiments in Sec.
+III show, that the generic chiral model applies for the magnetic field regions of the plateaus of quantized Hall resistance, while the Drude model describes the situation for all other magnetic fields.
+II.1 Coherent and chiral model of $n$-terminal device
+Our first model describes the limit of coherent and chiral current flow inside the 2DES of a sample equipped with $n$ ohmic contacts, cf. Fig. 2.
 Figure 2: Simplified circuit diagram of an $n$-terminal device.
@@
 circled arrow.
-Next, we derive a general model for the current in a 2DES device prepared in the coherent and chiral limit but equipped with $n$ ohmic contacts, cf.
-Fig.
-3.
-The model intends to describe current flow through a Hall bar in the region of a quantized plateau.
 The inner circle in Fig.
-3 indicates the coherent 2D conductor, the radial lines symbolize the $n$ terminals $i=1,2,...,n$.
-External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the coherent part of the sample and each contact.
-The current $I_{i}$ flows through the $i$th contact with ohmic resistance $R_{i}$ to ground.
-To describe the coherent inner part of the sample, we employ the Landauer-Büttiker formalism and write
+2 indicates the 2DES, the radial lines symbolize the $n$ terminals $i=1,2,...,n$ connected via ohmic contacts to the 2DES.
+External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the 2DES and each contact.
+The current $I_{i}$ flows through the $i$th contact via the ohmic resistance $R_{i}$ to ground.
+To describe the current flow through the 2DES (inner part of the sample), we employ the Landauer-Büttiker formalism (not to be confused with the LBP) and write
 $I_{i}=\sum_{j=1}^{n}G_{ij}V_{j}^{\prime}=(V_{i}^{\prime}-V_{i})/R_{i}\,,$ (2)
-where $G_{ij}$ are the coherent conductance coefficients, which can be formally derived from the scattering matrix of the problem.
-Assuming chiral transport, current flows only between adjacent contacts.
+where $G_{ij}$ are the conductance coefficients between mutual contacts.
+They can be formally derived from the scattering matrix of the problem.
+The left hand side of Eq.
+(2) is a general formulation of the transmission properties of the 2DES based on the Landauer-Büttiker formalism, while the last term of Eq.
+(2) applies Ohm’s law to the macroscopic contacts.
+If the transport within the 2DES is chiral, only those $G_{i,j\neq i}$ connecting adjacent contacts can be non-zero.
 The direction of the chirality is linked to the direction of the magnetic field vector.
-In our case as sketched in Fig.
-3, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}=1$ and $G_{ji}=0$ for $i=j+1$.
-The unitarity of the scattering matrix implies $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our chiral conductivity tensor takes the form
+In our case sketched in Fig.
+2, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}\neq 0$ for $i=j+1$.
+Assuming coherent transport, the scattering matrix is unitary, implying $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our coherent and chiral conductivity tensor takes the form
 $G_{ij}=\nu G_{0}\left\{\begin{matrix}1\,;&i=j+1\\[4.30554pt] -1\,;&i=j\hfill\par\LTX@newpage\\[4.30554pt] 0\,;&\text{else}\hfill\par\LTX@newpage\end{matrix}\right.\,,$ (3)
 where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels.
-The coherent limit applies for the quantized plateaus of the Hall resistance, where $\nu$ is an integer.
+While the assumption of coherence guaranties a unitary scattering matrix, for the validity of Eq.
+(3) it is actually sufficient to require the absence of momentum scattering.
+For the quantized plateaus of the Hall resistance $\nu$ is an integer.
 Inserting Eq.
 (3) into Eq.
-(5) and solving the equation system yields the currents
-$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)$ (7)
-where we introduced
+(2) yields a solvable linear equation system providing us with the prediction of the current into the $i$th contact
+$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)\,,$ (4)
 $P_{i}=\left\{\begin{matrix}\prod_{j=1}^{i}\left(1+\frac{1}{\nu G_{0}R_{j}}\right)\,;&i=1,2,...,n\\[4.30554pt] 1\,;&i=0\phantom{,2,...,n}\end{matrix}\right.$ (5)
@@
 $S_{i}=\sum_{j=1}^{i}P_{j-1}\frac{V_{j}}{\nu G_{0}R_{j}}\,.$ (6)
-Equation (7) describes the current distribution into the $n$ ohmic contacts of a chiral and coherent sample. For $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$ the currents can be written as
-$I_{i>1}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,I_{1}\,,$ (10)
-which are independent of $R_{1}$ as expected.
-II.3 Drude model of $n$-terminal device
-Our model describing coherent currents is based on the unitarity of the scattering matrix.
-However, regarding the actual macroscopic currents into various contacts, the unitarity condition implies the same result as we find from applying Kirchhoff’s laws.
-Therefore, if we replace $R_{\text{H}}$ by $(\nu G_{0})^{-1}$, for $n=3$ and $V_{2}=V_{3}=0$, Eq.
-(7) of the coherent and chiral model predicts the identical currents $I_{2,3}$ as the transversal current component $I_{xy}^{(2,3)}$ of the Drude model discussed in Sec.
-II.1 above.
-By extrapolating this observation, we can expand our prediction of the Drude model for the current distribution of an $n$-terminal junction and find
-$I_{i}$ $=$ $\frac{1}{1+(\mu B)^{2}}I_{xx}^{(i)}+\frac{(\mu B)^{2}}{1+(\mu B)^{2}}I_{xy}^{(i)}\quad\text{with}$
-$I_{xx}^{(i)}$ $=$ $\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (11)
-$I_{xy}^{(i)}$ $=$ $\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,.$
-The longitudinal component, $I_{xx}^{(i)}$, is determined by Ohm’s law while the transversal component, $I_{xy}^{(i)}$, is identical to Eq. (7), where we use $R_{\text{H}}=(\nu G_{0})^{-1}$.
-To correctly describe the regions of the plateaus of a quantized Hall resistance, we use constant and integer $\nu$ and assume $\mu=\infty$, a property of ballistic transport, such that $I_{xx}^{(i)}=0$.
-Then Eq.
-(II.3) is identical to Eq.
+A simplifying special case of some relevance is that of $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$, which yields the current ratios
+$\frac{I_{i>1}}{I_{1}}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,.$ (7)
+II.2 Drude model of an $n$-terminal device
+The Drude model provides a solution of Eq.
+(2), which describes the classical limit of diffusive and incoherent currents inside the 2DES.
+Using $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ we can rewrite Eq.
+(1) in terms of the magnetic-field-dependent conductivity tensor as
+$\begin{pmatrix}j_{x}\\ j_{y}\end{pmatrix}=\frac{\sigma_{0}}{1+(\mu B)^{2}}\begin{pmatrix}1&-\mu B\\ \mu B&1\end{pmatrix}\,\begin{pmatrix}E_{x}\\ E_{y}\end{pmatrix}\,,$ (8)
+where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$.
+For a laterally confined 2DES, in the steady state we expect $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
+The structure of the conductivity tensor allows us to partition the non-vanishing current density $j_{x}$ into two components with distinct physical meanings, namely
+$j_{x}\equiv j_{xx}+j_{xy}\quad\text{with}$
+$j_{xx}=\frac{\sigma_{0}}{1+(\mu B)^{2}}\,E_{x}\quad\text{and}\quad j_{xy}=\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{\text{H}}}\,E_{y}\,.$
+The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is the ohmic term, proportional to the specific conductivity and driven by the longitudinal electric field $E_{x}$.
+Caused by the additional Hall resistance at finite $B$, the ohmic current density is reduced by the factor $\left[1+(\mu B)^{2}\right]^{-1}$.
+The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is driven by the perpendicular electric field and attenuated by the non-local Hall resistance.
+For $\mu B\gg 1$ the ohmic component $j_{xx}$ vanishes and the current density $j\to j_{xy}$ becomes independent of $\sigma_{0}$.
+It implies, that momentum scattering is absent.
+At the same time, the transversal $j_{xy}$ is bent into a current carrying contact as it follows the charged edge of the Hall bar.
+Figure 3: Drude model: Simplifying sketch showing the partitioning of the current contributions $I_{xx}$ versus $I_{xy}$ at a three-terminal junction.
+A negative voltage is applied to the left contact, while the other contacts are grounded.
+$I_{xx}$ is devided according to Kirchhoff’s current rule.
+Neglecting the partial reflection of current at ohmic contacts, only one egde is negatively charged, other edges are at ground potential.
+$I_{xy}$ is guided by the charged edge and flows entirely from the left to the lower contact.
+In our experiments, we measure the current, which corresponds to the integral of the current densities in Eq.
+(II.2) across the Hall bar.
+For the partition of the current at a junction we can now use the following rules: The ohmic $I_{xx}$ branches according to Kirchhoffs current law.
+For simplicity, we neglect the small resistance $R_{0}$ of the 2DES ($R_{i}/R_{0}\sim 50$).
+The non-ohmic $I_{xy}$ partitions according to Eq.
+(4), if we include non-integer filling factors [$R_{\text{H}}=(\nu G_{0})^{-1}$].
+Therefore, in the limit $\mu B\gg 1$, the prediction for the current division of the Drude model becomes identical to that of our coherent and chiral model derived in Sec.
+II.1.
+To illustrate the different behaviors of $I_{xx}$ and $I_{xy}$, in Fig.
+3 we sketch the current division at a three-terminal junction.
+Combining both contributions, the Drude model predicts
+$I_{i}$ $=$ $I_{xx}^{(i)}+I_{xy}^{(i)}\quad\text{with}$
+$I_{xx}^{(i)}$ $=$ $\frac{1}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (10)
+$I_{xy}^{(i)}$ $=$ $\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,$
+for the current flowing into the $i$th contact.
+The original scope of the classical Drude model was a description of diffusive transport at $\mu B\lesssim 1$.
+Interestingly, in the limit $\mu B\to\infty$, the first term of Eq.
+(II.2) vanishes, $I_{xx}\to 0$, and the prediction of the Drude model becomes identical to Eq.
 (4).
+This property also means that for the limits $\mu B\lesssim 1$ versus $\mu B\gg 1$, Eq.
+(II.2) precisely describes both of our limiting models.
+For intermediate $\mu B$, Eq.
+(II.2) smoothly interpolates between the limits.
+This is useful, because $\mu(B)$ strongly oscillates due to the Shubnikov-de-Haas oscillations.
 III experimental results and discussion
-Multiterminal current measurements yield the distribution of the current into the different contacts.
-In this section, we compare the results of multi-terminal measurements and theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$.
-We will see, that multiterminal measurements can provide answers to the questions about chirality and the local current distribution in magnetotransport.
+Multiterminal current measurements yield the distribution of current into various contacts.
+In this section, we discuss the results of multi-terminal measurements and compare them with theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$.
+We will see, that even though multiterminal measurements do not provide the local current distribution across the Hall bar, their comparison with reasonable limiting models can provide answers to the questions about chirality of the current for the QHE.
 III.1 Sample and measurement setup
@@
 (a) Current ratios $-I_{3}/I_{1}$ and $-I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq.
-(II.3) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq.
-(II.3) with $\mu(B)$ as free parameter (dotted red line).
-(b) enlargements near $\nu=1$ showing experimental data and the Drude model prediction for $\mu=\mu_{0}$.
+(II.2) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq.
+(II.2) with $\mu(B)$ as free parameter (dotted red lines).
+(b) enlargements near $\nu=1$ showing experimental data, the Drude model prediction for $\mu=\mu_{0}$ and the model prediction of Eq.
+(4) for coherent and chiral current at $\nu=1$ (dotted red lines).
 Figure 6: Mobility $\mu(B)$ as determined from SdH-oscillations of the longitudinal resistance between contacts 2 and 3 (black dots) in comparison with $\mu(B)$ found by fitting Eq.
-(II.3) to the three-terminal measurements shown in Fig.
+(II.2) to the three-terminal measurements shown in Fig.
 5 (red squares).
-Figure 7: Four-terminal measurements for three different contact configurations, cf.
-insets, further configurations and labeling as in Fig.
-5.
-(a) symmetric four-terminal circuit.
-(b) and (c) variations of asymmetric four-terminal circuits.
+Figure 7: Four-terminal measurements for three different contact configurations; the blue numbers in the insets indicate grounded contacts, black numbers indicate floating contacts and the external voltage is applied to contact 1 (red); further details as in Fig.
+5; (a) symmetric four-terminal circuit; (b) and (c) asymmetric configurations.
 For multi-terminal current measurements, we apply the voltage of $V_{1}=\pm 1\,$mV to the source contact 1 (using a Keithley 2450 source meter).
@@
 As expected, we find that $|I_{3}|$ increases with $B$ while $|I_{5}|$ decreases accordingly, such that $I_{3}(B)+I_{5}(B)=-I_{1}(B)$.
-The dashed black line indicates the prediction of the Drude model according to Eq.
-(II.3) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$.
+The dashed black line indicates the prediction of the model according to Eq.
+(II.2) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$.
 For this large mobility, the Ohmic term (first term) of Eq.
-(II.3) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$).
-In this limit, with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$, Eq.
-(II.3) becomes identical to the coherent chiral model expressed in Eq.
-(7).
-The Drude model and the measured data agree in the classical limit for $B\lesssim 0.5$ T.
-As we increase $B$ further, the Drude model with a constant $\mu(B)=\mu_{0}$ still predicts the correct trend, but we observe growing deviations for filling factors away from integer values [top axis of Fig.
-5(a)].
-The enlargement around the first plateau of the QHE in Fig.
-5(b) clarifies, that model and experiment match exactly only at specific values of $B$ including integer bulk filling factors (here $\nu=1$).
-An exact match between measurements and model can be achieved by using $\mu(B)$ as a free fit parameter in Eq.
-(II.3).
+(II.2) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$), such that with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$ Eq.
+(4) [Drude model] and Eq.
+(II.2) [chiral model] become identical.
+In the classical limit for $B\lesssim 0.5$ T the Drude model agrees well with our measurement.
+As we increase $B$ further, the Drude model still predicts the correct trend but as long as we use a constant $\mu(B)=\mu_{0}$ (black dashed lines), we observe growing deviations.
+These are clearly visible in the enlargement around the first plateau at $\nu=1$ in Fig.
+5(b).
+The red dashed lines are the prediction of the chiral model according to Eq.
+(4) [or Eq.
+(II.2) using $\mu B=\infty$], where we inserted a constant filling factor $\nu=1$.
+It perfectly fits the plateau values of the current ratios.
+An exact match between measurements and model beyond the plateaus can be achieved by using $\mu(B)$ as a free fit parameter in Eq.
+(II.2).
 For the dotted red line in Fig.
@@
 [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to $T>4\,$K between the two measurements.
-The agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq.
-(II.3) to our three-terminal current measurements supports the validity of our model.
+The general agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq.
+(II.2) to our three-terminal current measurements supports the validity of our model.
 The two regions of divergent $\mu(B)$ in Fig.
-6 correspond to the plateau regions at $\nu=2$ and $\nu=1$.
-Here, $I_{xx}=0$ and the coherent and chiral model expressed in Eq.
-(7) is valid.
-Away from the plateau regions, we find strongly reduced mobility values with $201\lesssim\mu<$ 20, which indicate diffusive transport and a finite ohmic contribution $I_{xx}$ to the current.
-The overall agreement between the prediction of Eq.
-(II.3) and our three-terminal measurements is convincing.
-On the one hand, it supports the assumption of coherent and chiral transport for the quantized plateaus of the Hall resistance.
-On the other hand, it points to non-chiral transport described by the Drude model not only in the classical regime of the Hall effect but also for large $B$ in between the plateaus.
+6 correspond to the plateau regions with the local filling factors $\nu=2$ and $\nu=1$.
+Here, $I_{xx}$ vanishes.
+Away from the plateau regions, we find strongly reduced mobility values, indicating a transition to diffusive transport.
+The difference between the Drude model predictions for $\mu=\mu_{0}$ [dashed line in Fig.
+5(a)] and the measured currents is the ohmic current contribution $I_{xx}$, which becomes clearly visible for $\mu<20$.
+So-far we have discussed three-terminal current measurements.
+With the knowledge of the contact resistances of our sample, we can quantitatively compare the experiments with model predictions.
+Between the plateaus, the mobility is reduced and the Drude model fits perfectly to our data, indicating non-chiral transport.
+Regarding the plateau regions, our three-terminal measurements are clearly consistent with chiral transport.
+Importantly, any additional non-zero transmission coefficients in Eq.
+(3) due to non-chiral currents would modify the prediction of Eq.
+(4).
+Clearly, we do not observe such deviations.
 III.3 Four-terminal measurements
 In Fig.
-7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements.
+7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements (blue solid lines).
 Again, we measure the currents into contacts $i=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig.
@@
 If we assume $\mu(B)=\mu_{0}$, Eq.
-(II.3) of the Drude model reproduces the overall dependences $I_{i}(B)$ with a similar accuracy as it did for the three-terminal measurement (dashed black lines).
+(II.2) reproduces the overall dependences $I_{i}(B)$ with a similar accuracy and similar deviations as it did for the three-terminal measurement (dashed black lines).
 The dotted red lines correspond to the predictions of Eq.
-(II.3) using the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf.
+(II.2), where we used the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf.
 Fig.
 6.
-Again, for $B\lesssim 0.5\,$T, theory and experiment agree well.
-For $B>0.5\,$T, the agreement is still perfect for the plateau regions but becomes worse away from the plateaus of quantized Hall resistance.
-experimental results is even worse (for $B>0.5\,$T).
-This observation allows two immediate conclusions: (1) for the plateau regions, the chiral transport already found in the three-terminal measurement is confirmed; (2) away from the plateaus, the non-chiral transport is also confirmed.
-However, for more than three terminals, the model prediction of Eq.
-(II.3) fails for $B>0.5\,$T.
+As it was the case for the three terminal measurements, for the plateau regions, corresponding to diverging $\mu(B)$, the agreement between theory and the measured currents is perfect.
+However, away from the plateaus, the theory curves using the varying $\mu(B)$ does no longer yield a perfect agreement between theory and measured data.
+It means, that the division of the ohmic $I_{xx}$ component of the current is not accurately predicted.
+Nevertheless, our four-terminal measurements confirm the main conclusions of our three-terminal measurements: (1) Away from the plateaus, the transport is not chiral.
+(2) For the plateau regions, the agreement between our four-terminal measurements and the chiral and coherent model supports chiral transport.
+III.4 Limits of our model in the diffusive regime
+Away from the plateaus, the measured currents depend on the contact configuration beyond the predictions of our model.
+We conjecture, that our initial assumption of a homogeneous Hall bar fails near the intersections.
+In the regime of diffusive transport it implies a homogeneous electric field inside the Hall bar.
+In reality, near intersections, both components, $E_{x}$ and $E_{y}$ of the electric field and, therefore, the currents $I_{xx}$ and $I_{xy}$ become inhomogeneous.
+While $I_{xy}$ is still completely bent into the contacts, the division of $I_{xx}$ is influenced by its inhomogeneity.
+A more realistic model would require a suitable numerical calculation, e.g., based on a path integral formalism.
+We conjecture, that such a corrected model would yield perfect agreement with measurements, namely that the Drude model is still correct for more than three terminals, where it worked perfectly for three terminals.
+III.5 Purely experimental confirmation of non-chiral currents between plateaus
 In Fig. 8,
 Figure 8: Comparison of the respective largest measured current contributions flowing into a grounded contact for three different configurations of four-terminal measurements.
-The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled such that all three theory curves according to Eq.
-(II.3) with $\mu=\mu_{0}$ are identical following the edge between shaded and white backgrounds.
+The dotted horizontal lines depict the predicted current ratios for filling factors $1\leq\nu\leq 4$.
+The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled to account for the slightly different contact resistances, such that all three theory curves according to Eq.
+(II.2) with $\mu=\mu_{0}$ are identical and follow the edge between shaded and white backgrounds.
+Sketches: Arrows indicate the displayed current branches; colored numbers indicate current carrying and gray numbers floated contacts.
 we plot for each of the three four-terminal measurements shown in Fig.
-7 the branch corresponding to the largest current.
-We slightly scaled the data to compensate for small variations in the contact resistances.
-After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) and on the plateaus of the QHE.
-The edge between shaded and white backgrounds corresponds to the prediction of the Drude model for constant $\mu=\mu_{0}$ [dashed black lines in Fig.
+7 only the branch $I_{k}/I_{1}$ corresponding to the largest current.
+In a chiral model, this is the current flowing into the first grounded contact after emission from contact 1.
+We slightly scaled the data (based on the values provided in Table 1) to compensate for small variations in the contact resistances.
+After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) as well as on the plateaus of the QHE.
+The line linking shaded and white backgrounds corresponds to the prediction of Eq.
+(II.2) for constant $\mu=\mu_{0}$ [equal to the dashed black lines in Fig.
 7(a)].
-Compared to this model prediction, the current is reduced between the plateaus for all three geometries.
-The reduction is particularly strong for the set-up with contact 6 connected to ground.
-In the case of chiral transport, contact 6 is the last one to be reached by the current, while most current should flow into contact 3.
-The reduction of the current into contact 3 goes along with an enhancement of the current into contact 6, cf.
-Fig.
-7(b).
-This is a clear signature of non-chiral current but, the current reduction into contact 3 (and the current enhancement into contact 6) exceeds even the prediction of our model expressed by Eq.
-(II.3).
-Likewise, for the set-up with contact 2 grounded, the reduction of the current into contact 2 is smaller than the reduction predicted by Eq.
-(II.3).
-So, a smaller amount of the non-chiral current reaches contact 5 as it is much further away from contact 1 than contact 2.
-The details of the measured current distribution clearly depend on the contact configuration.
-This behavior demonstrates the limits of our model.
-For the derivation of Eq.
-(II.3), we assumed an infinitely long Hall bar without additional grounded contacts.
-Hence, for identical contact resistances, its prediction is independent of the variations between specific configurations of the four-terminal set-up (the resistance of the Hall bar itself at $B=0$ is much smaller than the contact resistances).
-To overcome the failure of our simplified model, it would be necessary to account for the more complicate steady-state electrostatics of a multi-terminal device, in particular the inhomogeneity of the electric field near a junction.
-Suitable numerical calculations could clarify this point.
-For the moment, we conjecture that such a corrected model would support our previous result based on the three-terminal measurement, namely that the Drude model correctly describes the transition regions between plateaus.
-IV Screening theory and chiral transport
-Figure 9: Coherent and chiral models for the first plateau of the QHE: Sketch of the distribution of the current $I_{xy}$ near $\nu=1$ for $V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{2}=R_{3}$ with $\vec{B}$ pointing into the plane of the 2DES.
-Color intensities indicate the distribution of the absolute value of the current density into the three branches.
-Black bent arrows show the direction and relative magnitudes of electron flow related with the charged edges indicated by blue minus signs for $V_{1}<0$.
-For $V_{1}>0$, current densities remain the same while the edge charges become positive and black arrows indicate the current flow.
-(a) Prediction of the Landauer Büttiker picture, assuming that current flows in one-dimensional chiral edge channels only.
-(b) Prediction of the screening theory for the case of edge ICSs (current flows inside ICSs only).
-(c) Prediction of the screening theory for the case of a bulk ICS.
-All three cases predict chiral current flow within the 2DES and result in identical currents $I_{1}$, $I_{2}$ and $I_{3}$ into the contacts.
-In the stem towards contact 1 we sketch the electrochemical potentials of the first Landau level (LL1, solid line) and the related Hall potential ($U_{\text{H}}$, dotted red line).
-In this section, we restrict the discussion to the plateau regions of the quantized Hall resistance, for which our multiterminal experiments indicate chiral and coherent transport.
-The unique characteristics of chiral transport of a 2D conductor is the opposite directions of current flow along the opposite edges of the conductor.
-The QHE reveals the additional feature of fundamental resistance values being fractions of $h/e^{2}$.
-These values can be measured with an extremely high accuracy and stability with deviations of less than $3\times 10^{-9}$ having been demonstrated between high resolution measurements of various samples in different laboratories [16].
-In order to explain the measured highly accurate resistance values in fractions of $h/e^{2}$, scattering has to be (almost) completely suppressed.
-The LBP [1, 2] is a single-particle model based on the assumption that the Landau levels form perfect 1D edge channels, which support chiral transport, while the bulk of the 2DES is assumed to be electrically insulating because of disorder-induced localization.
-The edge channels are compressible, meaning that there exist occupied and unoccupied states at the chemical potential.
-The mere assumption of scattering free transport via compressible edge channels as an explanation of the extremely accurate measurements of the fundamental resistance value $h/e^{2}$ is unsatisfactory.
-The screening theory offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$, which is based on first principle calculations [7, 8, 9].
-Already the earliest purely classical considerations, taking into account the Coulomb interaction, led to the realization of a segmentation of the 2DES confined in a Hall bar in perfectly screened compressible regions and completely unscreened ICSs [3, 4, 5, 6].
-Inside the ICSs, scattering is forbidden because the excitation of a carrier costs the energy $\hbar\omega_{\text{c}}$ of the Landau gap.
-While numerous experiments already support the screening theory [17, 10, 11, 12, 18, 19, 20, 21, 22, 23, 13], here our focus is on the question whether the screening theory is consistent with chiral transport.
-In Fig.
-9, we sketch the situation for the first plateau ($\nu=1$, $I_{xx}=0$) of the QHE for the case of a three-terminal junction.
-As a start, in Fig.
-9(a), we show the situation assumed in the LBP.
-All current flows inside the compressible edge channels (solid blue lines).
-The stem towards contact 1 contains a sketch of the first Landau level (LL1, solid line), which strictly follows the confinement potential for the single particle LBP.
-The edge channels emerge, where LL1 intersects with the chemical potential, which differs by $V_{\text{H}}$ between the respective edges.
-The Hall potential $U_{\text{H}}=eV_{\text{H}}$ (dotted line) drops linearly between the edge channels.
-To illustrate the screening theory [9], we consider the two limiting cases realized near the ends of the first plateau of the QHE.
-For simplicity, we neglect disorder effects.
-Near the lower magnetic field end with $\nu>1$, the screening theory predicts two narrow ICSs roughly resembling Landauer-Büttiker edge channels, cf.
-Fig.
-9(b).
-Near the high field end of the plateau, where $\nu\leq 1$, the prediction is a single bulk ICS, while the 2DES near the edges remains compressible, cf.
-Fig.
-9(c).
-Even for the first case with two edge ICSs, in contrast to the LBP, the current flows entirely inside the ICSs, where scattering is suppressed due to the Landau gap of the density of states.
-In the compressible regions of the 2DES, electrons can scatter and the local electric field is perfectly screened.
-As a result, the local electric field entirely drops inside the ICSs.
-The gradient of the Landau levels, i.e., the effective local electric field drives the current according to $\vec{j}=-en_{\text{s}}\,\vec{E}\times\vec{B}/B^{2}$.
-Therefore, no current flows in the perfectly screened compressible regions.
-The gradient of the potential energy of the Landau levels equals the gradient of the local electric field.
-As can be seen in the sketches of the first Landau level in Figures 9b and 9c, the gradient of the Landau levels has opposite directions inside the ICS(s) along the oposite edges of the sample.
-This is even the case if only a single ICS covers most of the 2DES, cf.
-Fig.
-9(c).
-Here, $\vec{E}=0$ and $\vec{j}=0$ in the center of the ICS and the current flows in opposite directions on both sides of the ICS.
-Consequently, even in case of a single bulk ICS, the current flow is still chiral and the local current density is larger near the sample edges, where the confinement potential drops steeply.
-V Summary and Outlook
-In summary, we have performed multi-terminal current measurements in the regime of the QHE and compared our results with predictions of a coherent and chiral model versus the Drude model.
-The Drude model perfectly describes our results in the classical regime of small magnetic fields, the coherent and chiral model fits for the regions corresponding to quantized plateaus of the Hall resistance.
-In the diffusive regions for large magnetic fields, i.e., between quantized plateaus our three-terminal experiment supports the predictions of the Drude model, which includes non-chiral transport.
-The four-terminal measurements demonstrate the limits of our model but clearly demonstrate non-chiral current contributions away from the plateaus of quantized Hall resistance.
-This work contributes to a deeper understanding of the microscopic dynamics and the distribution of current density inside realistic mesoscopic devices in the regime of the QHE.
+Compared to the model prediction, between the plateaus the current is reduced for all three geometries, which above we already interpreted in terms of a finite contribution of the non-chiral $I_{xx}$.
+Because the various contact resistances of our sample are almost identical, for a purely chiral current flow the branches of the largest current $I_{k}(B)/I_{1}$, shown in Fig.
+8 should be independent of the details of the contacts configuration (as long as the overall number of grounded contacts remains the same).
+The fact that $I_{k}(B)/I_{1}$ varies between configurations, is a direct experimental prove of non-chiral currents between the quantized plateaus.
+This conclusion experts additional support by a closer look at the amounts of current reduction: It systematically depends on the relative distances between contact 1 and the grounded contacts on both sample edges in a way that contradicts chiral current flow.
+III.6 Formulation of the proof of chiral current flow for quantized Hall resistance
+So far, we have proven non-chiral current flow between the quantized plateaus, while for the quantized plateaus we have merely stated that our experimental results are consistent with chiral current flow.
+However, our analysis allows a stronger statement, namely that the agreement between the chiral model and our measurements is a prove of chiral current flow.
+The proof is based on the coherent and chiral model introduced in Sec.
+II.1.
+Based on the Landauer-Büttiker formalism it describes the current division into $n$ contacts in terms of the mutual conductance coefficients $G_{ij}$ between the contacts.
+In this model, chiral current flow corresponds to the specific values of conductance coefficients given by Eq.
+(3).
+A non-chiral behavior would correspond to different values, including additional non-zero conductance coefficients in Eq.
+(3).
+These would inevitably result in different plateau values of the predicted current ratios.
+In all our measurements including three or four terminals shown in Fig.
+5, Fig.
+7 and Fig.
+8, we instead see a perfect agreement between our chiral model prediction and the measured plateau values of the current ratios.
+This is in particular evident when comparing the horizontal dotted lines in Fig.
+5(b) and Fig.
+8, which indicate the predictions of Eq.
+(4) for the corresponding integer filling factors, with the corresponding measured plateau values.
+This result stands in contrast to our finding in between plateaus, where the deviations of our measured currents from the prediction of Eq.
+(II.2) for $\mu B\gg 1$ provide a proof of non-chiral currents.
+IV Summary and Outlook
+We have performed multiterminal current measurements of a Hall bar in the regime of the QHE.
+We find that the current flow is chiral in the magnetic field regions of the quantized Hall resistance plateaus but becomes non-chiral in between the plateaus.
+Our result is based on the direct comparison of the measured current distribution into various contacts (three or four terminals) with two limiting models, which offer well understood, unique physical meanings.
+The classical Drude model serves us as the basis to describe non-chiral diffusive transport, while our second model is based on the generic chiral and coherent transmission problem.
+The comparison of different contact configurations for more than three terminals allows for a direct experimental proof of non-chiral currents, which confirm our model based results.
+Our work contributes to the understanding of the current flow inside mesoscopic devices in the regime of the QHE.
+It thereby confirms the assumption of chiral current flow frequently used for the case of quantized plateaus of the Hall resistance.
 Acknowledgement

### v3

@@
-Chiral nature of current flow in the regime of the quantized Hall effect
+Nature of current flow in the regime of the quantum Hall effect
 Abstract
@@
 It serves as a basis to understand the fractional, anomalous or spin QHEs, candidates for applications in quantum technology due to their topological properties.
-For optimizing all these applications it is essential to understand the current distribution inside the Hall bar, a question disputed for decades.
+For optimizing all these applications it is essential to understand where the current flows inside the Hall bar, a question disputed for decades.
 We perform multiterminal current measurements on a Hall bar and compare the results with limiting models.
-We show, based on these simple experiments, that the current flow is chiral for the plateaus of quantized Hall resistance.
-Everywhere else, the current is described by the Drude model.
+We confirm, based on these experiments, that the current flow is chiral for the plateaus of quantized Hall resistance.
+Everywhere else, including the ranges between plateaus, our results comply with the Drude model, which predicts homogeneous current flow across a homogeneous Hall-bar.
 I Introduction
@@
 This property gives rise to the chiral trajectory of a free charged particle in a homogeneous magnetic field, the mirror symmetry of the current in a two-terminal measurement, $I(\vec{B})=I(-\vec{B})$, or the point symmetry of the Hall voltage, $V_{\text{H}}(\vec{B})=-V_{\text{H}}(-\vec{B})$.
-In this article, we explore the implication of the axial symmetry of $\vec{B}$ for the current flow in the regime of the quantized Hall effect.
-In the classical equation of motion, $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$, the axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ acting on an electron with charge $-e$ and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to the fields $\vec{E}$ and $\vec{B}$.
-The solution for a ballistic electron is an in-plane spiral-shaped trajectory composed of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}\neq\vec{\upsilon}$ and a cyclotron motion with radius $R_{\text{c}}=|\vec{\upsilon}|/\omega_{\text{c}}$, where $\omega_{\text{c}}=eB/m$ is the cyclotron frequency, cf.
-Fig.
-1(a).
-If we define the chirality of a trajectory in analogy to the handedness of a screw thread, the above trajectory of a ballistic electron is left-handed (and would be right handed for a positive charge).
-Figure 1: Motion (expressed by drift velocity) of an electron confined to the $xy$-plane with in-plane electric and perpendicular magnetic fields, $\vec{E}$ and $\vec{B}$ (red arrows), respectively.
-(a) Ballistic limit (free electron).
-The spiral is the classical trajectory consisting of a uniform drift velocity $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ and a (left-handed) cyclotron motion with its radius being proportional to the initial velocity.
-(b) Diffusive limit.
-For $|\mu B|<1$ the spiral motion is damped.
-However, the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ and tilts it by the Hall angle $\varphi$.
-(c) Diffusive Hall bar in the steady state limit with edges parallel to the applied $\vec{E}$ (red arrow).
-The Hall voltage compensates the Lorentz force, $-e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$ (or $E_{y}=|\vec{F}_{\text{L}}|/e$), such that the drift velocity is $\vec{\upsilon}_{\text{d}}$ as for $\vec{B}=0$.
-The Drude model describes magnetotransport in the classical limit of the diffusive regime.
-Its key parameter is the electron mobility $\mu$.
-For $\mu B\lesssim 1$ (corresponding to a mean-free path $l_{\text{m}}\lesssim R_{\text{c}}$), the cyclotron motion is impeded by Markovian momentum scattering.
-Consequently, the Drude model neglects the cyclotron motion by assuming $\vec{\upsilon}=\vec{\upsilon}_{\text{d}}$.
-Applying the relaxation ansatz $\dot{\vec{\upsilon}}_{\text{d}}=\vec{\upsilon}_{\text{d}}/\tau$ with the momentum scattering time $\tau=m\mu/e$, the solution of the equation of motion introduced above in the diffusive limit becomes
-$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left(\vec{E}-\vec{E}\times\mu\vec{B}\right)\,\text{ for }\vec{B}\perp\vec{E}\,.$ (1)
-[For $\vec{B}\not\perp\vec{E}$ there would be a third summand inside the parantheses, namely $\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}$.] For finite $B$, $\vec{j}$ is bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$ in respect to $\vec{E}$, cf.
+In the regime of the quantum Hall effect (QHE), the situation is more complicated because of an inhomogeneous distribution of the electric field in case of a quantized Hall resistance.
+As a result, the current density distribution in the quantum Hall regime has been a matter of discussion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] since the discovery of the QHE [21].
+In this article, we use multiterminal measurements to explore the implication of the axial symmetry of $\vec{B}$ for the current flow in the regime of the QHE.
+Multiterminal measurements are not capable of directly detecting the current density distribution, but based on a careful analysis they can nevertheless reveal the nature of current flow, e.g., if it is chiral or not.
+Before we discuss our experimental results, we introduce relevant theoretical aspects and models for comparison with our measurements.
+In the classical equation of motion (EOM), $m\dot{}\vec{\upsilon}=-e\vec{E}+\vec{F}_{\text{L}}$, the axial character of $\vec{B}$ is expressed in terms of the vector product of the Lorentz force $\vec{F}_{\text{L}}=-e\vec{\upsilon}\times\vec{B}$ acting on an electron with charge e-e and effective mass $m$ at the momentary velocity $\vec{\upsilon}$, exposed to a magnetic field and an electric field $\vec{E}$.
+[…]
+Figure 1: Motion (expressed by drift velocity) of an electron confined to the xyxy-plane with in-plane electric and perpendicular magnetic field, $\vec{E}\perp\vec{B}$ (red arrows).
+(a) Ballistic limit (free electron): The spiral indicates the cyclotron motion of an electron consisting of a uniform drift velocity $\vec{\upsilon}_{\text{d}}$ and a (left-handed) circular motion with its radius being proportional to the initial velocity.
+(b) Diffusive limit: the magnetic field reduces the drift velocity by the factor $1/\sqrt{(1+(\mu B)^{2}}$ compared to its absolute value for $B=0$ and tilts it by the Hall angle $\varphi$.
+The circular motion of the free electron is damped.
+(c) Diffusive Hall bar with edges parallel to the applied $\vec{E}$: The Hall voltage compensates the Lorentz force, $e\vec{\nabla}V_{\text{H}}=\vec{F}_{\text{L}}$, the drift velocity $\vec{\upsilon}_{\text{d}}$ is parallel to the sample edges.
+Next, we consider the diffusive regime, which is dominated by momentum scattering.
+In its classical limit, its electron dynamics can be described by the Drude model with the electron mobility $\mu$ as central parameter.
+For $\mu B\lesssim 1$ (corresponding to a mean-free path $l_{\text{m}}\lesssim R_{\text{c}}$), momentum scattering counteracts the circular motion.
+Consequently, the Drude model omits the circular motion and replaces the velocity in the EOM (above) by the uniform drift velocity $\dot{}\vec{\upsilon}\to\dot{}\vec{\upsilon}_{\text{d}}$.
+After applying the relaxation time ansatz $\dot{\vec{\upsilon}}_{\text{d}}=-\vec{\upsilon}_{\text{d}}/\tau$, where $\tau=m\mu/e$ is the momentum scattering time, the EOM becomes a linear vector equation.
+Its solution,
+$\vec{\upsilon}_{\text{d}}=\frac{-\mu}{1+(\mu B)^{2}}\,\left[\vec{E}-\vec{E}\times\mu\vec{B}+\mu^{2}(\vec{E}\cdot\vec{B})\vec{B}\right]\,$, (1)
+can be simply determined by rearranging the EOM.
+For $\vec{B}\perp\vec{E}$ the term proportional to $\vec{E}\cdot\vec{B}$ vanishes and $\vec{\upsilon}_{\text{d}}$ is confined to a plane perpendicular to $\vec{B}$.
+Compared to $\vec{\upsilon}_{\text{d}}=-\mu\vec{E}$ for $B=0$, a finite magnetic field reduces its absolute value according to $|\vec{\upsilon}_{\text{d}}|=\mu E/\sqrt{1+(\mu B)^{2}}$.
+In respect to the direction of $\vec{E}$, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, is then bent by the Hall angle $\varphi=\sphericalangle(\vec{j},\vec{E})=\tan^{-1}(\mu B)$, cf.
 Fig.
 1(b).
-Importantly, as long as both the carrier density $n_{\text{s}}$ and the mobility $\mu$ are homogeneous, the current density $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$ is homogeneous, too.
-While the Drude model is designed to describe the limit of $\mu B\lesssim 1$, its prediction for the drift velocity for $\mu B\to\infty$ where the friction forces in Eq.
-(1) become negligible, namely $\vec{\upsilon}_{\text{d}}\to\vec{E}\times\vec{B}/B^{2}$, is identical to the ballistic solution.
-Since for $\mu B\to\infty$ the rapid cyclotron motion averges out over reasonable length scales, both models predict identical current densities, too, for $\mu B\to\infty$.
-In the remainder of the article, we restrict ourselves to a two-dimensional electron system (2DES), which defines the $xy$-plane, with an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular homogeneous magnetic field $\vec{B}=(0,0,B)$.
+Importantly, as long as both the carrier density, $n_{\text{s}}$, and the mobility, $\mu$, are homogeneous, $\vec{j}$ is homogeneous, too.
+As argued above, the Drude model applies for $\mu B\lesssim 1$.
+Interestingly, for the opposite limit, namely $\mu B\to\infty$ and $R_{\text{c}}\ll l_{\text{m}}$, the rapid circular motion of the cyclotron trajectory is not damped but averages out over reasonable length scales.
+At the same time, for $\mu B\to\infty$, the friction forces in Eq.
+(1) become negligible, such that the current density predicted by the Drude model becomes $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}\to-en_{\text{s}}\vec{E}\times\vec{B}/B^{2}$, which is independent of $\mu$ and identical to the solution for ballistic electrons.
+In the remainder of the article, we restrict ourselves to a degenerate two-dimensional electron system (2DES), which defines the $xy$-plane, with an in-plane electric field $\vec{E}=(E_{x},E_{y},0)$ and a perpendicular homogeneous magnetic field $\vec{B}=(0,0,B)$.
 This is a common scenario realized in many mesoscopic magnetotransport measurements.
@@
 In the steady state, we expect $j_{y}=0$ everywhere even for $B\neq 0$, because the edges reflect electrons.
-Applied to Eq.
-(1) $j_{y}=0$ yields $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$ with the Hall resistance defined as $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$.
-Physically, the transversal field $E_{y}$ is sourced by charge accumulated at the sample edges.
-With $E_{y}=\vec{\nabla}V_{\text{H}}$ it also gives rise to the Hall voltage $V_{\text{H}}$, that can be measured across the Hall bar.
-Because in steady state, the Lorentz force, acting on the carriers, is exactly canceled by $-eE_{y}$ (the forces associated with $B$ and $E_{y}$ cancel each other), inside the Hall bar the electrons move as if only the longitudinal electric field $E_{x}$ was applied.
-Given a homogeneous sample (homogeneous distributions of $n_{\text{s}}$ and $\mu$ away from the edges of the Hall bar), the current density $j_{x}$ is uniform and the current is $I=j_{x}b$, with $b$ being the width of the Hall bar.
-While the Drude model predicts a homogeneous current density for the classical Hall effect, the axial character of the magnetic field and the breaking of the time-reversal symmetry are manifest in the edge charges and the related Hall voltage.
-In addition, the Hall resistance is non-local as it does not cause any dissipation inside the Hall bar: At $B=0$ the two terminal current measurement of a Hall bar is the sum of the ohmic sample resistance and the resistance of the leads (including the ohmic contacts), $R_{\text{2T}}(0)=R_{0}+R_{\textrm{leads}}$.
-At finite $B$, the non-local Hall resistance enhances the resistance of the Hall bar, such that $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}+R_{\textrm{leads}}$.
-As we increase $B$, at low enough temperatures the quantization of the density of states into discrete Landau-levels results in the quantum Hall effect (QHE).
-Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance ($R_{0}=0$).
-The latter indicates the absence of momentum scattering within the Hall bar for a quantized Hall resistance, in agreement with the classical Drude prediction for $\mu B\to\infty$.
+Applying $j_{y}=0$ to Eq.
+(1) yields the classical Hall effect with $E_{y}=-\mu BE_{x}$ and $j_{x}=E_{y}/R_{\text{H}}$, where we introduced the Hall resistance $R_{\text{H}}=-\frac{B}{n_{\text{s}}e}$.
+The transversal field $E_{y}$ is generated by charge accumulated at the sample edges, which can be measured in terms of the Hall voltage $V_{\text{H}}$, related to $E_{y}$ via $E_{y}=|\vec{\nabla}V_{\text{H}}|$.
+Because in steady state, the Lorentz force, acting on the carriers, is exactly canceled by $eE_{y}$, the Drude model predicts that inside a Hall bar the electrons move as if only the longitudinal electric field $E_{x}$ was applied.
+Given a homogeneous sample (homogeneous distributions of $n_{\text{s}}$ and $\mu$ away from the edges of the Hall bar), the current density $j_{x}$ is then predicted to be uniform and the current becomes $I=j_{x}b$, with $b$ being the width of the Hall bar.
+Although the Drude model predicts a homogeneous current density for the classical Hall effect, the axial character of the magnetic field is actually manifest in the edge charges and the related Hall voltage, which break the time-reversal symmetry.
+Interestingly, other than an ohmic resistance, the Hall resistance does not add dissipation inside the Hall bar.
+(Remember, the related transversal field compensates the Lorentz force.) Instead, the Hall resistance is non-local.
+It causes additional dissipation where the current flow becomes parallel to $\vec{\nabla}V_{\text{H}}$, which typically happens near or inside the current carrying contacts.
+(For the quantum Hall regime, an enhanced dissipation near the contacts was experimentally confirmed [22, 23]).
+For a constant current, the overall dissipation is proportional to the two-terminal resistance, $R_{\text{2T}}$, of the Hall bar sample.
+At $B=0$, it is the sum of the resistances of the Hall bar and the leads (including the ohmic contacts), $R_{\text{2T}}(0)=R_{0}+R_{\textrm{leads}}$.
+At finite $B$, the resistance can be described as $R_{\text{2T}}(B)=\sqrt{R_{0}^{2}+R_{\text{H}}^{2}}+R_{\textrm{leads}}$.
+As we increase $B$, at low enough temperatures the quantization of the density of states into discrete Landau-levels results in the quantum Hall effect.
+Its main features are plateaus of quantized Hall resistance accompanied by a vanishing longitudinal resistance ($R_{0}\to 0$).
+The latter indicates the absence of momentum scattering within the Hall bar for a quantized Hall resistance.
+The related Shubnikov-de-Haas (SdH) oscillations of the longitudinal resistance reveal in the regime of the QHE an oscillating $R_{0}(B)$ with growing maxima between the quantized plateaus as $B$ is increased, cf.
+Fig.
+13 in Appendix C.
 To explain the quantized values of $R_{\text{H}}$, the Landauer-Büttiker picture (LBP) assumes the formation of chiral, one-dimensional (1D) edge channels, such that backward scattering is suppressed [2, 3].
-The conductance of a spin-resolved 1D channel without momentum scattering is $e^{2}/h$, which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$.
+The conductance of a spin-resolved 1D channel without momentum scattering is $e^{2}/h$ [24], which concurs with the observed quantization of the Hall resistance, $R_{\text{H}}=R_{\text{K}}/\nu$, in fractions of the von-Klitzing constant $R_{\text{K}}=h/e^{2}$.
 The filling factor $\nu$ describes the number of edge channels and is equal to two-times the filling fraction of the (spin-degenerate) Landau levels in the 2D bulk.
-The LBP assumes that the edge channels form where the potential energy of the Landau levels, which increase towards the depleted sample edges, intersect with the Fermi level [1, 2].
-To obtain finite ranges of quantized Hall resistances, additionally the bulk (between the edge channels) is assumed to be insulating due to Anderson localization.
-The direction of the current flow in the edge channels is predetermined by the gradient of the confinement potential.
-This yields chiral current flow, more precisely, current flow in opposite directions along opposite edges.
-The LBP provides an intuitive explanation for the quantized Hall resistance and the vanishing longitudinal resistance of the QHE.
-While its merit is its simplicity, its predictions are limited due to the phenomenological character as well as its neglect of interactions and non-equilibrium effects.
-In particular, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [3, 4].
-Hence, neglecting the Coulomb interaction between carriers in the LBP causes the formation of highly charged stripes between the edge channels.
-This local polarization corresponds to an unrealistically large Coulomb energy.
-Another weak point of the LBP is the coexistence of occupied and unoccupied states inside the current carrying edge channels at the chemical potential.
-Even, if we assume chiral edge channels, an almost complete suppression of carrier scattering, as indicated by the accuracy of the measurements of the von Klitzing constant, is unlikely under any realistic conditions.
-The screening theory [3, 4, 5, 6, 7, 8, 9] offers an alternative explanation of the highly accurate resistance values in fractions of $h/e^{2}$.
+The LBP assumes that the edge channels form where the Landau levels, whose potential energies increase towards the depleted sample edges, intersect with the Fermi level [2, 3].
+To obtain finite ranges of quantized Hall resistances, additionally the bulk (between the edge channels) is assumed to be insulating due to localization in a disorder potential in a high magnetic field [3].
+The direction of the current flow in the 1D edge channels is predetermined by the gradient of the confinement potential.
+This yields current flow in opposite directions along opposite edges.
+Following the usual practice, we call this chiral current flow.
+Depending on the direction of the magnetic field the chiral current flow is clockwise or counter-clockwise.
+The LBP provides an intuitive explanation for the quantization of the Hall resistance and the accompanying vanishing longitudinal resistance of the QHE.
+While its merit is its simplicity, it neglects interactions and non-equilibrium effects, which gives it a phenomenological character with a limited range of validity.
+For example, the single particle assumption of the LBP leads to the prediction of a stepwise increase of the carrier density at the edges of the Hall bar because of the successive population of Landau levels [5].
+It means, that neglecting the Coulomb interaction between carriers in the LBP causes the formation of highly charged stripes between the edge channels.
+This local polarization would give rise to an unrealistically large Coulomb energy.
+Another problem is the coexistence of occupied and unoccupied states at the chemical potential inside the current carrying edge channels according to the LBP.
+Even, if we take into account chiral current flow, an almost complete suppression of carrier scattering, as indicated by the high accuracy of the measurements of the von Klitzing constant [25], would then be unlikely under realistic conditions.
+The screening theory [5, 26, 27, 28, 29, 30, 14] offers an alternative explanation of the QHE.
 By taking into account the direct Coulomb interaction between electrons, it overcomes the limitations of the LBP.
-The Coulomb interaction causes a rearrangement of the free carriers such that the highly charged stripes are avoided.
-Early purely classical formulations considered this screening based on the Thomas-Fermi approximation [3, 4].
-They demonstrated that the Landau-quantization leads to a segmentation of the 2DES of a Hall bar in compressible regions and incompressible strips (ICSs).
+It is based on the realization, that the Coulomb interaction of free carriers allows screening of the local electric field by means of rearrangement of carriers.
+This screening helps to avoid the highly charged stripes and thereby lower the free energy.
+However, the screening is hindered, if Landau-quantization in a strong magnetic field causes an energy gap.
+Early purely classical formulations included screening based on the Thomas-Fermi approximation [5, 26].
+By minimizing the free energy, they demonstrated that the Landau-quantization leads to a segmentation of the 2DES of a Hall bar in compressible regions and incompressible strips (ICSs).
 The term ICS denotes areas of the 2DES, where the chemical potential lies within a local energy gap between completely filled and empty Landau levels.
 For this reason, at low enough temperature, inside the ICSs carriers cannot scatter and screening of the electric field is absent.
-ICSs are accompanied by adjacent compressible regions, in which a partly filled Landau-level is aligned with the chemical potential.
+ICSs are accompanied by adjacent compressible regions, in which a partly filled Landau-level is exactly aligned with the chemical potential.
 Hence, inside compressible regions, carriers can scatter.
-Provided the vicinity of unscreened ICSs, the screening of the local electric field becomes perfect inside the compressible regions.
+Provided the vicinity of unscreened ICSs, the screening of the local electric field becomes perfect inside the compressible regions [5].
 The early calculations already show, that both, width and position of the ICSs depend on the magnetic field as well as the local electron density distribution $n_{\text{s}}(x,y)$ at $B=0$.
-Later, for given boundary conditions the number and geometry of the ICSs were predicted more accurately using a self-consistent numerical calculation of Poisson’s equation within the Thomas-Fermi approximation and accounting for the quantum mechanical wavefunctions of the electrons in a mean field approach [7, 8, 9] (later confirmed in a full self-consistent approach of the quantum- electrostatic problem [10]).
-The calculations predict, that towards the higher magnetic field end of each plateau the ICSs merge into a single extended incompressible region centered in the Hall bar [7, 8, 9].
-These predictions clearly confirmed the electrostatic landscape measured in scanning probe spectroscopy experiments as a function of the magnetic field [11, 12, 13].
+Later, for given boundary conditions, the number and geometry of the ICSs were predicted more accurately using a self-consistent numerical calculation of Poisson’s equation within the Thomas-Fermi approximation and accounting for the quantum mechanical wavefunctions of the electrons in a mean field approach [29, 30, 14] (later confirmed in a fully self-consistent approach of the quantum-electrostatic problem [31]).
+The calculations predict, that towards the higher magnetic field end of each plateau the ICSs merge into a single extended incompressible region (bulk ICS) centered in the Hall bar [29, 30, 14].
+These calculations confirmed the electrostatic landscape measured in high-resolution scanning probe spectroscopy experiments as a function of the magnetic field [12, 32, 33], which where indicated earlier by scanning electro-optical (Pockels) effect experiments in larger samples [4, 34].
 According to the screening theory the electric field vanishes inside the perfectly screened compressible regions while it is exclusively contained inside the ICSs.
-For suppressed scattering the drift velocity of the electrons is $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$.
-Hence, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, vanishes in the compressible regions, while all current flows inside the ICSs [4, 14, 9, 15].
-In case of a bulk ICS, current should then flow in the center of the Hall bar.
-Recently, we confirmed the predicted bulk current directly by means of an additional tiny ohmic contact in the center of a Hall bar [16].
-Both, the electrostatics and the existence of bulk currents, predicted by the screening theory [9] and confirmed in measurements [11, 12, 13, 16] are in direct contrast to the assumptions of the LBP.
-Therefore, it is important to find out, whether another assumption of the LBP, which is often used, remains correct, namely whether the current flow is chiral for the quantized plateaus of the Hall resistance.
-Even within the screening theory this is still a controversal question [3, 4, 9, 10, 15].
-In the present article, we explore the question of chirality experimentally by employing multiterminal current measurements.
-Thereby, we probe the division of the current into various ohmic contacts of the Hall bar.
-Clearly, we do not directly probe the local current density distribution.
-Nevertheless, it is possible to distinguish between chiral versus non-chiral transport as the current division into multiple contacts directly depends on it.
-To quantify this dependence, we use the properties of the macroscopic scattering matrix, which yields the division of the current into various contacts as follows: In the case of chiral current flow, only unidirectional transmission coefficients connecting adjacent contacts are non-zero.
-In comparison, non-chiral current flow requires additional non-zero transmission coefficients, which would ultimately yield a different divison of the current into the various contacts.
-Here, we show that the measured current division can be described for arbitrary magnetic fields by either one of two limiting models, introduced in Sec.
-II below.
+For suppressed scattering ($\mu B\gg 1$) the drift velocity of the electrons is $\vec{\upsilon}_{\text{d}}=\vec{E}\times\vec{B}/B^{2}$ (ballistic electrons).
+Hence, the current density, $\vec{j}=-en_{\text{s}}\vec{\upsilon}_{\text{d}}$, vanishes in the compressible regions, while all current flows inside the ICSs [26, 35, 14].
+In case of a bulk ICS, current should then flow near the center of the Hall bar, while at the same time the Hall resistance should remain quantized.
+Recently, we precisely confirmed this predicted bulk current by exploring current flow through an additional tiny ohmic contact in the center of a Hall bar [36].
+Both, the electrostatics and the existence of bulk currents, predicted by the screening theory [14] and confirmed in measurements [4, 34, 12, 32, 33, 36] are in direct contrast to the assumptions of the LBP.
+Therefore, it is important to find out, whether another assumption of the LBP, based on the axial symmetry of the magnetic field vector, remains correct, namely whether the current flow is chiral for the quantized plateaus of the Hall resistance and whether the chirality survives even in the case of bulk current.
+The distribution of the local current density in the quantum Hall regime has been controversally discussed for some time, in particular, in articles going beyond the LBP [5, 26, 9, 8, 35, 14, 17, 20, 31, 36].
+In the present article, we present experiments designed to explore the question of the chirality of the current flow by employing multiterminal current measurements.
+In detail, we probe the division of the current into various ohmic contacts of the Hall bar.
+Measuring currents into contacts, we cannot directly probe the local current density distribution.
+Nevertheless, it is possible to distinguish between chiral versus non-chiral transport as the current division into multiple contacts directly depends on it, see Sec.
+III.3 for related measurements.
+To quantify the current distribution into various contacts, we use the properties of the macroscopic scattering matrix, which yields the division of the current into various contacts as follows: In the case of chiral current flow, only unidirectional transmission coefficients connecting adjacent contacts are non-zero.
+In comparison, non-chiral current flow requires additional non-zero transmission coefficients.
+Clearly, the two cases lead to different distributions of the current into the various contacts.
 In Sec.
-III, we then present our three- and four-terminal current experiments, compare them with the predictions of the two limiting models and discuss our results.
+II we introduce the two limiting models, the first one describing coherent and chiral current flow and the second one describing diffusive current (Drude model for multiple terminals).
+In Sec.
+III, we then present our three- and four-terminal current experiments and show that the measured current distribution into various contacts can be described for arbitrary magnetic fields by either one of two limiting models.
+Our results suggest, that the current flow is chiral for the quantized plateaus but unidirectional between the plateaus.
 II models
 Our first model generically assumes coherent and chiral current flow.
-The condition of coherence is technical.
-In practice, it is sufficient to assume the absence of scattering.
+The condition of coherence is assumed for convenience.
+While this assumption is technically not necessary, the absence of scattering in an ICS would actually imply coherent current flow.
 The second model applies the classical Drude model to a multiterminal current measurement.
-It describes diffusive and non-chiral current.
-Our experiments in Sec.
-III show, that the generic chiral model applies for the magnetic field regions of the plateaus of quantized Hall resistance, while the Drude model describes the situation for all other magnetic fields.
+It describes incoherent, diffusive and non-chiral current.
 II.1 Coherent and chiral model of n-terminal device
@@
 The inner circle indicates coherent and chiral current flow inside the 2DES.
-Each contact $i=1,...,n$ has an ohmic resistance $R_{i}$ and a voltage $V_{i}$ applied in respect to the electrical ground.
-$V_{i}^{\prime}$ are the voltages resulting at the intersections between the ohmic contacts and the coherent inner circle.
+Each contact k=1,,nk=1,...,n has an ohmic resistance $R_{k}$ and a voltage $V_{k}$ applied in respect to the electrical ground.
+$V_{k}^{\prime}$ are the voltages resulting at the intersections between the ohmic contacts and the coherent inner circle.
 The magnetic field vector pointing into the plane of the 2DES gives rise to a right handed chirality of a free charge which translates into a left handed chirality of the device, cf.
@@
 The inner circle in Fig.
-2 indicates the 2DES, the radial lines symbolize the $n$ terminals $i=1,2,...,n$ connected via ohmic contacts to the 2DES.
-External voltages $V_{i}$ can be applied to the contacts in respect to ground, while $V_{i}^{\prime}=V_{i}+I_{i}R_{i}$ are the voltages present at the intersects of the 2DES and each contact.
-The current $I_{i}$ flows through the $i$th contact via the ohmic resistance $R_{i}$ to ground.
+2 indicates the 2DES, the radial lines symbolize the $n$ terminals k=1,2,,nk=1,2,...,n connected via ohmic contacts to the 2DES.
+External voltages $V_{k}$ can be applied to the contacts in respect to ground, while $V_{k}^{\prime}=V_{k}+I_{k}R_{k}$ are the voltages present at the intersects of the 2DES and each contact.
+The current $I_{k}$ flows through the $k$th contact via the ohmic resistance $R_{k}$ to ground.
 To describe the current flow through the 2DES (inner part of the sample), we employ the Landauer-Büttiker formalism (not to be confused with the LBP) and write
-$I_{i}=\sum_{j=1}^{n}G_{ij}V_{j}^{\prime}=(V_{i}^{\prime}-V_{i})/R_{i}\,,$ (2)
-where $G_{ij}$ are the conductance coefficients between mutual contacts.
+$I_{k}=\sum_{j=1}^{n}G_{kj}V_{j}^{\prime}=(V_{k}^{\prime}-V_{k})/R_{k}\,$, (2)
+where $G_{kj}$ are the conductance coefficients between mutual contacts.
 They can be formally derived from the scattering matrix of the problem.
-The left hand side of Eq.
+The left-hand side of Eq.
 (2) is a general formulation of the transmission properties of the 2DES based on the Landauer-Büttiker formalism, while the last term of Eq.
 (2) applies Ohm’s law to the macroscopic contacts.
-If the transport within the 2DES is chiral, only those $G_{i,j\neq i}$ connecting adjacent contacts can be non-zero.
+If the transport within the 2DES is chiral, only those $G_{k,j\neq k}$ connecting adjacent contacts can be non-zero.
 The direction of the chirality is linked to the direction of the magnetic field vector.
 In our case sketched in Fig.
-2, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{ij}\neq 0$ for $i=j+1$.
-Assuming coherent transport, the scattering matrix is unitary, implying $\sum_{i}G_{ij}=\sum_{j}G_{ij}=0$, such that our coherent and chiral conductivity tensor takes the form
-$G_{ij}=\nu G_{0}\left\{\begin{matrix}1\,;&i=j+1\\[4.30554pt] -1\,;&i=j\hfill\par\LTX@newpage\\[4.30554pt] 0\,;&\text{else}\hfill\par\LTX@newpage\end{matrix}\right.\,,$ (3)
+2, we consider $\vec{B}$ pointing into the plane of the 2DES, such that $G_{kj}\neq 0$ for $k=j+1$.
+Assuming coherent transport for the inner circle of Fig.
+2, the scattering matrix is unitary, implying $\sum_{k}G_{kj}=\sum_{j}G_{kj}=0$, such that our coherent and chiral conductivity tensor takes the form
+[…]
 where $G_{0}=e^{2}/h$ and $\nu$ is the filling factor of the (spin-resolved) Landau-levels.
-While the assumption of coherence guaranties a unitary scattering matrix, for the validity of Eq.
-(3) it is actually sufficient to require the absence of momentum scattering.
-For the quantized plateaus of the Hall resistance $\nu$ is an integer.
+The assumption of coherence, which guarantees unitarity of the scattering matrix and the validity of Eq.
+(3) is expected for the absence of scattering inside an ICS.
+For the quantized plateaus of the Hall resistance, $\nu$ is an integer.
+The Eq.
+(3) might be considered as our formal definition of chiral current flow.
 Inserting Eq.
 (3) into Eq.
-(2) yields a solvable linear equation system providing us with the prediction of the current into the $i$th contact
-$I_{i}=\frac{V_{i}^{\prime}-V_{i}}{R_{i}}\quad\text{with}\quad V_{i}^{\prime}=\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)\,,$ (4)
-$P_{i}=\left\{\begin{matrix}\prod_{j=1}^{i}\left(1+\frac{1}{\nu G_{0}R_{j}}\right)\,;&i=1,2,...,n\\[4.30554pt] 1\,;&i=0\phantom{,2,...,n}\end{matrix}\right.$ (5)
-and
-$S_{i}=\sum_{j=1}^{i}P_{j-1}\frac{V_{j}}{\nu G_{0}R_{j}}\,.$ (6)
-A simplifying special case of some relevance is that of $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$, which yields the current ratios
-$\frac{I_{i>1}}{I_{1}}=\frac{{P_{n}}/{P_{i}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{i}}\,.$ (7)
+(2) yields a solvable linear equation system providing us with the prediction of the current into the $k$th contact
+$I_{k}=\frac{1}{R_{k}}\left[\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)-V_{k}\right]\,$, (4)
+where we used $V_{k}^{\prime}=\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)$ with
+$P_{k}=\left\{\begin{matrix}\prod_{j=1}^{k}\left(1+\frac{1}{\nu G_{0}R_{j}}\right)\,;&k=1,2,...,n\\[4.30554pt] 1\,;&k=0\phantom{,2,...,n}\end{matrix}\right$. (5)
+$\text{and }S_{k}=\sum_{j=1}^{k}P_{j-1}\frac{V_{j}}{\nu G_{0}R_{j}}\,.$ (6)
+In Appendix A we provide a derivation of Eq. (4). A simplifying special case of some relevance is that of $V_{1}\neq 0$ but all other applied voltages $V_{i>1}=0$, which yields the current ratios
+$\frac{I_{k>1}}{I_{1}}=\frac{{P_{n}}/{P_{k}}}{1-{P_{n}}/{P_{1}}}\,\frac{1}{\nu G_{0}R_{k}}\,$ (7)
+irrespective of the sign of $V_{1}$. To visualize the nature of chiral current flow, in Fig. 3
+Figure 3: Chiral current flow in a y-junction: Simplifying sketches for a three-terminal junction for $\vec{B}$ pointing into the plane.
+In panel (a) a negative, in (b) a positive voltage $V_{1}$ is applied to contact 1, while the other contacts are grounded.
+Contact resistances are neglected, such that charge accumulates at the edge between contacts 1 and 2, only.
+Due to the direction of the applied magnetic field current flows clockwise (electrons counter clockwise).
+Bright blue arrows indicate the path way of the current generated by $V_{1}$.
+For $V_{1}<0$ it flows from contact 2 to 1; for $V_{1}>0$ it flows from 1 to 3 and then to 2.
+The faint arrows indicate a persistent chiral current that flows still for $V_{1}=V_{2}=V_{3}=0$.
+we sketch the chiral current path predicted by our model for just three terminals with $V_{2}=V_{3}=0$, while we neglect the contact resistances for simplicity.
+Arrows indicate the technical current direction, the electrons flow in the opposite direction.
+The clockwise chiral current flow is fixed by the direction of $\vec{B}$ pointing from the surface downwards through the 2DES plane.
+Contacts 2 and 3 are grounded, while $V_{1}<0$ in panel (a) and $V_{1}>0$ in (b).
+The chiral current predicted by Eq.
+(3) includes a persistent contribution carried by states below the chemical potentials of all three leads, which flows even for $V_{1}=V_{2}=V_{3}=0$.
+Such a persistent current is expected for the plateaus of the QHE [26, 35].
+The faint arrows indicate persistent current, while the bright blue arrows indicate the flow of the sum of the persistent current and the additional current imposed by applying $V_{1}\neq 0$.
+For $V_{1}<0$, excess electrons move directly from contact 1 to contact 2 ($1\to 2$) according to $V_{1}<V_{2}=V_{3}=0$, hence, the imposed current flows from $2\to 1$.
+For $V_{1}>0$, the number of electrons flowing from $2\to 3$ and that from $3\to 1$ are equal, while the number of electrons flowing from $1\to 2$ is smaller.
+Therefore, the imposed current flows from $1\to 3\to 2$.
+In summary, the path of the chiral current is fixed by the direction of the magnetic field.
+As a result, if we neglect contact resistances, all current generated by applying a voltage to contact 1, always flows through contact 2 independent of the sign of $V_{1}$ or the number of additional grounded contacts.
+Upon adding contact resistances, Eq.
+(4) predicts corrections, such as some current flowing through additional contacts.
 II.2 Drude model of an n-terminal device
@@
 where the specific conductivity (defined at $B=0$) is $\sigma_{0}=en_{\text{s}}\mu$.
-For a laterally confined 2DES, in the steady state we expect $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
+For a 2DES laterally confined in the $y$-direction, in the steady state we expect $j_{y}=0$, which implies $E_{y}=-\mu BE_{x}$ and, consequently, $j_{x}=\sigma_{0}E_{x}=R_{\text{H}}^{-1}E_{y}$.
 The structure of the conductivity tensor allows us to partition the non-vanishing current density $j_{x}$ into two components with distinct physical meanings, namely
@@
 The longitudinal component $j_{xx}=\sigma_{xx}E_{x}$ is the ohmic term, proportional to the specific conductivity and driven by the longitudinal electric field $E_{x}$.
-Caused by the additional Hall resistance at finite $B$, the ohmic current density is reduced by the factor $\left[1+(\mu B)^{2}\right]^{-1}$.
-The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is driven by the perpendicular electric field and attenuated by the non-local Hall resistance.
-For $\mu B\gg 1$ the ohmic component $j_{xx}$ vanishes and the current density $j\to j_{xy}$ becomes independent of $\sigma_{0}$.
-It implies, that momentum scattering is absent.
-At the same time, the transversal $j_{xy}$ is bent into a current carrying contact as it follows the charged edge of the Hall bar.
-Figure 3: Drude model: Simplifying sketch showing the partitioning of the current contributions $I_{xx}$ versus $I_{xy}$ at a three-terminal junction.
-A negative voltage is applied to the left contact, while the other contacts are grounded.
-$I_{xx}$ is devided according to Kirchhoff’s current rule.
-Neglecting the partial reflection of current at ohmic contacts, only one egde is negatively charged, other edges are at ground potential.
-$I_{xy}$ is guided by the charged edge and flows entirely from the left to the lower contact.
-In our experiments, we measure the current, which corresponds to the integral of the current densities in Eq.
-(II.2) across the Hall bar.
-For the partition of the current at a junction we can now use the following rules: The ohmic $I_{xx}$ branches according to Kirchhoffs current law.
-For simplicity, we neglect the small resistance $R_{0}$ of the 2DES ($R_{i}/R_{0}\sim 50$).
-The non-ohmic $I_{xy}$ partitions according to Eq.
-(4), if we include non-integer filling factors [$R_{\text{H}}=(\nu G_{0})^{-1}$].
-Therefore, in the limit $\mu B\gg 1$, the prediction for the current division of the Drude model becomes identical to that of our coherent and chiral model derived in Sec.
+Caused by the additional Hall resistance, at finite $B$ the ohmic current density is reduced by the factor $\left[1+(\mu B)^{2}\right]^{-1}$ and vanishes for $\mu B\gg 1$.
+The transversal component $j_{xy}=\sigma_{xy}E_{y}$ is proportional to the perpendicular electric field.
+For $\mu B\gg 1$, with $j_{xy}\to E_{y}/R_{\text{H}}$ it becomes independent of the ohmic resistance.
+This implies, that for $\mu B\gg 1$ momentum scattering is absent.
+At the same time, the transversal $j_{xy}$ follows the charged edge of the Hall bar.
+This way, it is bent into a current carrying contact.
+In our experiments, we measure the current, which corresponds to the integral across the Hall bar of the current densities in Eq.
+(II.2).
+To illustrate the contributions of $I_{xx}$ and $I_{xy}$, in Fig.
+4
+Figure 4: Diffusive current flow in a y-junction (Drude model): Simplifying sketches showing the partitioning of the current contributions $I_{xx}$ versus $I_{xy}$ in a three-terminal junction for $\vec{B}$ pointing into the plane.
+In panel (a) a negative, in (b) a positive voltage is applied to contact 1, while the other contacts are grounded.
+Contact resistances are neglected, such that charge accumulates at the edge between contacts 1 and 2, only.
+$I_{xx}$ is devided according to Kirchhoff’s current law.
+$I_{xy}$ is guided by the charged edge and flows perpendicular to the Hall field $E_{y}$ (indicated by dashed arrows).
+$I_{xy}$ flows from contact 2 to 1 (opposite to electron flow).
+we sketch the current division according to the Drude model in a three-terminal junction for the identical conditions as in Fig.
+3, where we sketched current path ways within the chiral model, ($V_{1}\neq 0$, $V_{2}=V_{3}=0$, $R_{1}=R_{2}=R_{3}=0$).
+$I_{xx}$ branches according to Kirchhoff’s current law, while the non-ohmic $I_{xy}$ is bent by the edge charges (respective magnetic field) into the next available (non-floating) contact in current direction.
+The classical Drude model predicts unidirectional current flow, which is in no way chiral.
+Nevertheless, the current component $I_{xy}$ is bent such that it flows between contacts 1 and 2, as does the imposed current predicted by our chiral model in Sec.
 II.1.
-To illustrate the different behaviors of $I_{xx}$ and $I_{xy}$, in Fig.
-3 we sketch the current division at a three-terminal junction.
-Combining both contributions, the Drude model predicts
-$I_{i}$ $=$ $I_{xx}^{(i)}+I_{xy}^{(i)}\quad\text{with}$
-$I_{xx}^{(i)}$ $=$ $\frac{1}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{\sum_{m=1}^{n}V_{m}/R_{m}}{\sum_{m=1}^{n}1/R_{m}}-V_{i}\right]$ (10)
-$I_{xy}^{(i)}$ $=$ $\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{i}}\left[\frac{1}{P_{i}}\left(\frac{S_{n}}{P_{n}-1}+S_{i}\right)-V_{i}\right]\,$
-for the current flowing into the $i$th contact.
+Applying the Drude model to our general $n$-terminal sample allows us to predict the currents $I_{k}$ flowing into each contact in the diffusive regime.
+If we assume a homogeneous 2DES in our Hall bar and neglect possible asymmetries in the contact configuration, we may incorporate the resistance $R_{0}(B)$ of the 2DES into the $n$ contact resistances, and find
+$I_{k}$ $=$ $I_{xx}^{(k)}+I_{xy}^{(k)}\quad\text{with}$
+$I_{xx}^{(k)}$ $=$ $\frac{1}{1+(\mu B)^{2}}\,\frac{1}{R_{k}}\left[\frac{\sum_{j=1}^{n}V_{j}/R_{j}}{\sum_{j=1}^{n}1/R_{j}}-V_{k}\right]$ (10)
+$I_{xy}^{(k)}$ $=$ $\frac{(\mu B)^{2}}{1+(\mu B)^{2}}\,\frac{1}{R_{k}}\left[\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)-V_{k}\right]\,$
+for the current flowing into the $k$th contact.
+In Appendix Sec.
+B we provide a detailed explanation of this result and discuss the relevance of $R_{0}(B)$ for the case of an asymmetric configuration of contacts.
 The original scope of the classical Drude model was a description of diffusive transport at $\mu B\lesssim 1$.
-Interestingly, in the limit $\mu B\to\infty$, the first term of Eq.
+However, in the limit $\mu B\to\infty$, the first term of Eq.
 (II.2) vanishes, $I_{xx}\to 0$, and the prediction of the Drude model becomes identical to Eq.
-(4).
-This property also means that for the limits $\mu B\lesssim 1$ versus $\mu B\gg 1$, Eq.
-(II.2) precisely describes both of our limiting models.
+(4), see Appendix B for a discussion.
+Therefore, we can interpret Eq.
+(II.2) as follows: for the limit $\mu B\lesssim 1$, which occurs for non-integer $\nu$ between the plateaus of the QHE, Eq.
+(II.2) describes diffusive transport based on the Drude model.
+For the opposite limit $\mu B\gg 1$, which occurs for integer $\nu$ on the plateaus, Eq.
+(II.2) precisely describes transport according to our chiral model.
 For intermediate $\mu B$, Eq.
 (II.2) smoothly interpolates between the limits.
-This is useful, because $\mu(B)$ strongly oscillates due to the Shubnikov-de-Haas oscillations.
+This is useful, because $\mu(B)$ strongly oscillates due to the SdH oscillations.
 III experimental results and discussion
 Multiterminal current measurements yield the distribution of current into various contacts.
-In this section, we discuss the results of multi-terminal measurements and compare them with theory predictions in a wide magnetic field region corresponding to $1\lesssim\nu<\infty$.
-We will see, that even though multiterminal measurements do not provide the local current distribution across the Hall bar, their comparison with reasonable limiting models can provide answers to the questions about chirality of the current for the QHE.
+In this section, we discuss the results of such experiments and compare them with theory predictions in a wide magnetic field range corresponding to filling factors between $1\lesssim\nu<\infty$.
 III.1 Sample and measurement setup
 We use a Hall bar as sketched in Fig. 5.
-Figure 4: Sketch of a Hall bar with six contacts, characterized by their ohmic resistances, $R_{i}$.
-We apply a voltage $V_{1}$ to contact 1 and measure currents into contacts 1, 3, and 5.
-Contacts 2, 4, and 6 can be left floating or connected to electrical ground.
-Current amplifiers at contacts 3 and 5 result in additional voltages of $\simeq 20\mu$V applied to contacts 3 and 5 (not shown, but taken into account for generating model curves).
-$\vec{B}$ points into the plane of the 2DES.
-It is etched from a GaAs / (Al,Ga)As heterostructure, which contains a two-dimensional electron system (2DES) 130 nm beneath the surface.
-We performed the experiments at a temperature of $T\simeq 300\,$mK in a He-3 evaporation cryostat.
-At this temperature and zero magnetic field, the electron density and mobility of the 2DES, determined from Hall measurements, are $n_{\textrm{s}}\simeq 1.2\times 10^{11}\textrm{ cm}^{-2}$ and $\mu\simeq 3.95\times 10^{6}\textrm{cm}^{2}/\textrm{Vs}$ (yielding a mean-free-path of $\lambda_{\text{m}}\simeq 23\,\mu$m), respectively.
+[…]
+[…]
+For multi-terminal current measurements, we apply the voltage of $V_{1}=\pm 1\,$mV to the source contact 1 (using a Keithley 2450 source meter).
+We measure the current $I_{1}(B)$ flowing from the Hall bar into contact 1, using the source meter, and the currents $I_{3}(B)$ and $I_{5}(B)$ flowing into contacts 3 or 5, respectively, using Basel Precision Instr.
+current amplifiers.
+Contacts 3 and 5 are connected to the electrical ground, while the remaining contacts 2, 4, and 6 may be either electrically floating or connected to ground depending on the experiment.
+For a first impression, in Fig.
+6
+Figure 6: Currents $I_{k}(B)$ measurement of a symmetric y-junction: Contacts 2, 4 and 6 in Fig.
+5 are floating, $V_{1}=+1\,$mV for $B>0$ and $V_{1}=-1\,$mV for $B<0$, contacts 3 and 5 are connected to the electrical ground.
+Measured currents are $I_{1}$ (black line), $I_{3}$ (red) and $I_{5}$ (blue), gray lines correspond to $I_{1}$.
+We find $I_{3}>I_{5}$ for $B>0$ and $I_{5}>I_{3}$ for $B<0$, where $I_{1}+I_{3}+I_{5}=0$.
+we present the measured currents $I_{1}(B)$, $I_{3}(B)$ and $I_{5}(B)$ for two three-terminal measurements (contacts 2, 4 and 6 are floating, such that contacts 3 and 5 correspond to contacts 2 and 3 in the sketches in Fig.
+3 and Fig.
+4).
+We define $B>0$ for $\vec{B}$ pointing downwards (from the surface through the 2DES) and $B<0$ for $\vec{B}$ pointing upwards.
+We applied the voltages $V_{1}=+1\,$mV and swept $B>0$ versus $V_{1}=-1\,$mV for $B<0$.
+In both cases, for $B=0$, the currents $I_{3}\simeq I_{5}$, because the contact resistances are almost equal.
+(Of course, the sign of the currents depends on the sign of the voltage.) As expected, $I_{1}=I_{3}+I_{5}$ always $|I_{1}|$ decreases as $|B|$ is increased because of $R_{\text{H}}\propto|B|$.
+Importantly, for $B>0$, we observe $|I_{3}|>|I_{5}|$, while for $B<0$, we find $|I_{5}|>|I_{3}|$.
+This behavior is independent of the sign of $V_{1}$ (not shown here).
+These data directly visualize our prediction that the current distribution in the various grounded contacts is governed by the direction of $\vec{B}$, only, cf.
+Eq.
+(II.2) or the sketches in Fig.
+3 and Fig.
+4.
+[…]
+contact $k$ resistance $R_{k}$ ($\Omega$) voltage $V_{k}$ (mV) geometry $a_{k}$
+1 2420 1.000 1
+2 2420 0.5
+3 2400 -0.021 1
+4 2350 1
+5 2420 -0.018 1
+6 2400 0.5
+Table 1: Parameters of the circuit sketched in Fig.
+5.
+Contact resistances $R_{k}$ include ohmic contacts, cables and RC-filters.
+Voltage $V_{1}=1$ mV is applied to contact 1, $V_{3}$ and $V_{5}$ are offset voltages of the current amplifiers connected to the respective contacts.
+Contacts 2, 4 and 6 are either floating or directly connected to ground.
+Hence, $V_{2}$, $V_{4}$ and $V_{6}$ are zero (variable), if the respective contact is grounded (floated).
+To account for the contribution of the Hall bar itself to the resistance, we defined a geometry parameter and use $R_{k}\to R_{k}+a_{k}/[n_{\text{s}}e\mu(B)]$ in the model calculations of the ohmic contributions; $a_{2}$ and $a_{6}$ are smaller, as they are closest to contact 1, such accounting for a shorter current path that comes into effect when the respective contacts are grounded.
+The correction is negligible for $B=0$.
+This preparation allows us to perform quantitative comparisons between measurements and models.
+For the reminder of the article we discuss experiments with $V_{1}=1\,$mV and $B>0$.
+For the chiral model, electrons then move from contact to contact along the counterclockwise pathway $1\to 2\to 3\to 4\to 5\to 6\to 1$.
+Since the external voltage that we apply to contact 1 exceeds all other voltages by far (and all contact resistances are similar), for a sizable magnetic field $B>0$, according to the prediction of both models and our first experiment shown in Fig.
+6, the largest current will flow between contacts 1 and the next grounded contact following in counterclockwise direction.
+III.2 Three-terminal measurements
+We start our quantitative comparison between measurements and models for the three-terminal measurements already presented in Fig.
+6 above.
+Here, the grounded contacts 3 and 5 are arranged in a mirror symmetric geometry in respect to the horizontal axis through contact 1.
+In Fig.
+7(a),
 Figure 7: Three-terminal measurement with symmetric circuit, cf.
 inset in (a).
-A voltage of $V=1\,$mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating.
+A voltage of V=1V=1\,mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating.
 $\vec{B}$ points from the surface into the 2DES plain such that $I_{xy}$ is bent towards contacts 2 and 3.
-(a) Current ratios $-I_{3}/I_{1}$ and $-I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq.
+(a) Current ratios $I_{3}/I_{1}$ and $I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq.
 (II.2) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq.
@@
 Figure 9: Four-terminal measurements for three different contact configurations; the blue numbers in the insets indicate grounded contacts, black numbers indicate floating contacts and the external voltage is applied to contact 1 (red); further details as in Fig.
-5; (a) symmetric four-terminal circuit; (b) and (c) asymmetric configurations.
-For multi-terminal current measurements, we apply the voltage of $V_{1}=\pm 1\,$mV to the source contact 1 (using a Keithley 2450 source meter).
-We measure the current $I_{1}(B)$ flowing from the Hall bar into contact 1, using the source meter, and the currents $I_{3}(B)$ and $I_{5}(B)$ flowing into contacts 3 or 5, respectively, using Basel Precision Instr.
-current amplifiers.
-Contacts 3 and 5 are connected to the electrical ground, while the remaining contacts 2, 4, and 6 may be either electrically floating or connected to ground depending on the experiment, cf.
-Fig.
-4.
-Here, we apply $B$ perpendicular to the 2DES, such that it points from the sample surface towards the 2DES, hence, electrons contributing to $I_{xy}$ move from contact to contact along the counterclockwise pathway $1\to 2\to 3\to 4\to 5\to 6\to 1$.
-Due to the direction of the magnetic field and a voltage applied to contact 1, the largest portion of $I_{xy}$ will always flow between contact 1 and contact 3, if contact 2 is floating, or to contact 2, if it is grounded.
-Whether this current is positive or negative depends on the sign of $V_{1}$.
-We performed a series of measurements at $B=0$ to determine the contact resistances ($R_{i}\sim 2.5\,\text{k}\Omega$) and offset voltages ($V_{3,5}\sim 20\,\mu$V) of the two current amplifiers.
-In addition, we used the comparison between measured data and model curves for a fine tuning of these circuit parameters, which remain constant between measurements.
-They are listed in Table 1.
-contact $i$ resistance $R_{i}$ ($\Omega$) voltage $V_{i}$ (mV) geometry $a_{i}$
-1 2420 1.000 1
-2 2420 0 0.5
-3 2400 -0.021 1
-4 2350 0 1
-5 2420 -0.018 1
-6 2400 0 0.5
-Table 1: Parameters of the circuit sketched in Fig.
-4.
-Contact resistances $R_{i}$ include ohmic contacts, cables and $RC$-filters.
-Voltage $V_{1}$ is applied to contact 1, $V_{2}$ and $V_{3}$ are offset voltages of the current amplifiers.
-To account for the contribution of the Hall bar itself to the resistance, we defined a geometry parameter and use $R_{i}\to R_{i}+a_{i}/[n_{\text{s}}e\mu(B)]$ in the model calculations of the ohmic contributions.
-The correction is negligible for $B=0$.
-III.2 Three-terminal measurements
-First, we discuss a three-terminal measurement in a mirror symmetric geometry, for which we keep the contacts 2, 4 and 6 electrically floating; current flows from contact 1 into contacts 3 and 5.
-In Fig.
-5(a), we present as solid (blue) lines the measured current ratios $-I_{3}(B)/I_{1}(B)$ and $-I_{5}(B)/I_{1}(B)$.
-As expected, we find that $|I_{3}|$ increases with $B$ while $|I_{5}|$ decreases accordingly, such that $I_{3}(B)+I_{5}(B)=-I_{1}(B)$.
-The dashed black line indicates the prediction of the model according to Eq.
-(II.2) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$.
-For this large mobility, the Ohmic term (first term) of Eq.
-(II.2) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$), such that with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$ Eq.
-(4) [Drude model] and Eq.
-(II.2) [chiral model] become identical.
+7; (a) symmetric four-terminal circuit; (b) and (c) asymmetric configurations.
+we present as solid (blue) lines the measured current ratios $I_{3}(B)/I_{1}(B)$ and $I_{5}(B)/I_{1}(B)$.
+As already discussed above, $I_{3}/I_{1}$ increases with $B$ while $I_{5}/I_{1}$ decreases accordingly, such that $I_{3}(B)/I_{1}(B)-I_{5}(B)/I_{1}(B)=1$.
+The dashed black line indicates the prediction of the Drude model according to Eq.
+(II.2) for $n=3$, if we use the constant mobility $\mu(B)=\mu_{0}=395\,\textrm{m}^{2}/\textrm{Vs}$ that we measured at $B=0$.
 In the classical limit for $B\lesssim 0.5$ T the Drude model agrees well with our measurement.
@@
 These are clearly visible in the enlargement around the first plateau at $\nu=1$ in Fig.
-5(b).
-The red dashed lines are the prediction of the chiral model according to Eq.
-(4) [or Eq.
-(II.2) using $\mu B=\infty$], where we inserted a constant filling factor $\nu=1$.
-It perfectly fits the plateau values of the current ratios.
+7(b).
+The red dotted lines in Fig.
+7(b) are the prediction of the chiral model according to Eq.
+(4) [corresponding to Eq.
+(II.2) for $\mu B=\infty$] and a constant filling factor $\nu=1$.
+It perfectly fits the values of the current ratios along the $\nu=1$ plateau of the quantized Hall resistance.
 An exact match between measurements and model beyond the plateaus can be achieved by using $\mu(B)$ as a free fit parameter in Eq.
@@
 For the dotted red line in Fig.
-5(a), we used the magnetic field dependent values of $\mu(B)$ as shown by red squares in Fig.
-6.
-For comparison, we have also determined $\mu(B)$ from the Shubnikov-de-Hass oscillations of a standard longitudinal resistance measurement, see the black dots in Fig.
-6.
-[^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to $T>4\,$K between the two measurements.
-The general agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq.
-(II.2) to our three-terminal current measurements supports the validity of our model.
+7(a), we used the magnetic field dependent values of $\mu(B)$, which are plotted as red squares in Fig.
+8.
+For comparison, we have also determined $\mu(B)$ from the SdH oscillations of a standard longitudinal resistance measurement, see the black dots in Fig.
+8  [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to T>4T>4\,K between the two measurements..
+The general agreement between the $\mu(B)$ values obtained from two completely independent measurements, namely longitudinal resistance measurements versus fitting Eq.
+(II.2) to our three-terminal current measurements, strongly supports the validity of our model.
 The two regions of divergent $\mu(B)$ in Fig.
 8 correspond to the plateau regions with the local filling factors $\nu=2$ and $\nu=1$.
-Here, $I_{xx}$ vanishes.
-Away from the plateau regions, we find strongly reduced mobility values, indicating a transition to diffusive transport.
+Here, $I_{xx}$ vanishes supporting $\mu B\to\infty$.
+Away from the plateau regions, in Fig.
+8 we find mobility values strongly reduced even compared to $\mu_{0}$.
+This indicates a transition from ballistic transport on the plateaus to diffusive transport between the plateaus.
 The difference between the Drude model predictions for $\mu=\mu_{0}$ [dashed line in Fig.
-5(a)] and the measured currents is the ohmic current contribution $I_{xx}$, which becomes clearly visible for $\mu<20$.
+7(a)] and the measured currents [blue line] is precisely the ohmic current contribution $I_{xx}$, which becomes clearly detectable for $\mu B<20$.
+To understand the meaning of the perfect agreement of the Drude model fits based on a variable $\mu(B)$ it is useful to mention, that for a large enough mobility, here $\mu B\gtrsim 20$, the Ohmic term (first term) of Eq.
+(II.2) becomes too small to be detected.
+Neglecting $I_{xx}$ in Eq.
+(II.2) and using $(\nu G_{0})^{-1}=R_{\text{H}}(B)$, then makes Eq.
+(4) [chiral model] and Eq.
+(II.2) [Drude model] mathematically identical.
 So-far we have discussed three-terminal current measurements.
 With the knowledge of the contact resistances of our sample, we can quantitatively compare the experiments with model predictions.
-Between the plateaus, the mobility is reduced and the Drude model fits perfectly to our data, indicating non-chiral transport.
-Regarding the plateau regions, our three-terminal measurements are clearly consistent with chiral transport.
+Between the plateaus, the mobility is strongly reduced and the Drude model fits perfectly to our data, indicating non-chiral diffusive transport.
+Regarding the quantized plateau regions, our three-terminal measurements are clearly consistent with Eq.
+(4), describing chiral current flow, if we assume a constant integer filling factor (for the entire plateau).
+It confirms, that for the quantized Hall plateaus the current flows within regions of constant integer filling factor, while the bulk filling factor (averaged over the width of the Hall bar) continuously decreases proportional to $1/B$.
 Importantly, any additional non-zero transmission coefficients in Eq.
@@
 (4).
-Clearly, we do not observe such deviations.
+Clearly, the fact that Eq.
+(4) perfectly fits our data for the plateaus with $\mu B\to\infty$ (without the need of additional non-zero transmission coefficients) suggests chiral transport for the quantized Hall plateaus.
 III.3 Four-terminal measurements
+Three-terminal measurements are special, because if a model curve agrees with one current branch, say $I_{3}/I_{1}$ in Fig.
+7, because of $I_{1}+I_{3}+I_{5}=0$ it will also agree with the second branch, here $I_{5}/I_{1}$.
+Clearly, this is not the case if we add additional terminals.
+Then, any inaccuracies of our model will be visible as deviations between measurements and model predictions.
+In that sense, experiments with four or more terminals provide a stricter test of our models compared to the three-terminal measurements above.
 In Fig.
-7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements (blue solid lines).
-Again, we measure the currents into contacts $i=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig.
-7(a), contact 6 in (b), and contact 2 in (c).
+9, we present current ratios $I_{k}/I_{1}$ for three different four-terminal measurements (blue solid lines).
+Again, we measure the currents into contacts $k=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig.
+9(a), contact 6 in (b), and contact 2 in (c).
 If we assume $\mu(B)=\mu_{0}$, Eq.
-(II.2) reproduces the overall dependences $I_{i}(B)$ with a similar accuracy and similar deviations as it did for the three-terminal measurement (dashed black lines).
-The dotted red lines correspond to the predictions of Eq.
-(II.2), where we used the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf.
-Fig.
-6.
+(II.2) reproduces the overall dependences $I_{k}(B)$ with a similar accuracy and similar deviations as it did for the three-terminal measurement (dashed black lines).
+The dotted red lines correspond to model curves of Eq.
+(II.2) using the $\mu(B)$ plotted in Fig.
+8 and determined by fitting the three-terminal measurement shown in Fig.
+7.
 As it was the case for the three terminal measurements, for the plateau regions, corresponding to diverging $\mu(B)$, the agreement between theory and the measured currents is perfect.
-However, away from the plateaus, the theory curves using the varying $\mu(B)$ does no longer yield a perfect agreement between theory and measured data.
-It means, that the division of the ohmic $I_{xx}$ component of the current is not accurately predicted.
-Nevertheless, our four-terminal measurements confirm the main conclusions of our three-terminal measurements: (1) Away from the plateaus, the transport is not chiral.
+This observation supports our previous conclusion, that in the plateau regions the current flow is chiral, while away from the plateaus, the chiral model fails.
+Noticeable, away from the plateaus, where the current flow is diffusive, the agreement between theory curves and measured data is no longer perfect.
+Thereby, that the best agreement occurs for the symmetric contact configuration in Fig.
+9(a).
+This behavior suggests, that our model is inaccurate in the diffusive transport regime.
+Thereby, the deviations depend on the details of the contact configuration.
+In conclusion, our four-terminal measurements confirm the main findings of our three-terminal measurements: (1) Away from the plateaus, the transport is not chiral.
 (2) For the plateau regions, the agreement between our four-terminal measurements and the chiral and coherent model supports chiral transport.
 III.4 Limits of our model in the diffusive regime
-Away from the plateaus, the measured currents depend on the contact configuration beyond the predictions of our model.
-We conjecture, that our initial assumption of a homogeneous Hall bar fails near the intersections.
-In the regime of diffusive transport it implies a homogeneous electric field inside the Hall bar.
-In reality, near intersections, both components, $E_{x}$ and $E_{y}$ of the electric field and, therefore, the currents $I_{xx}$ and $I_{xy}$ become inhomogeneous.
-While $I_{xy}$ is still completely bent into the contacts, the division of $I_{xx}$ is influenced by its inhomogeneity.
-A more realistic model would require a suitable numerical calculation, e.g., based on a path integral formalism.
-We conjecture, that such a corrected model would yield perfect agreement with measurements, namely that the Drude model is still correct for more than three terminals, where it worked perfectly for three terminals.
-III.5 Purely experimental confirmation of non-chiral currents between plateaus
+The deviations between the prediction of Eq.
+(II.2) and our measurements in the diffusive regime point to an incomplete model for $I_{xx}$.
+The problem is not related with the different contact resistances and offset voltages of the current amplifiers, because they are taken into account in Eq.
+(II.2).
+Instead, it is related with the mesoscopic dynamics of the electrons inside the Hall bar.
+In Sec.
+II.2 we determined the distribution of $I_{xx}$ into various contacts based on the contact resistances and the Hall resistance in series, while treating the Hall bar itself as a current node.
+To do this, we assumed a homogeneous current density distribution across the Hall bar.
+However, in particular near the intersections between the Hall bar and its contacts the distributions of the electric field and the current density become locally inhomogeneous.
+We conjecture, that such local inhomogeneities near the intersections cause the observed deviations between measured data and model curves in Fig.
+9.
+Note that the inhomogeneities affect the distribution of $I_{xx}$ into various contacts, while $I_{xy}$ is nevertheless bent into the contacts due to the charge accumulated along the edge of the Hall bar.
+A more realistic model would require a suitable numerical calculation, e.g., based on a path integral formalism with realistic boundary conditions.
+III.5 Experimental confirmation of non-chiral currents between plateaus
 In Fig. 10,
@@
 The dotted horizontal lines depict the predicted current ratios for filling factors $1\leq\nu\leq 4$.
-The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled to account for the slightly different contact resistances, such that all three theory curves according to Eq.
+The current ratios $I_{k}\,/\,I_{1}$ marked with a * are slightly scaled to account for the slightly different contact resistances, such that all three theory curves according to Eq.
 (II.2) with $\mu=\mu_{0}$ are identical and follow the edge between shaded and white backgrounds.
@@
 We slightly scaled the data (based on the values provided in Table 1) to compensate for small variations in the contact resistances.
-After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) as well as on the plateaus of the QHE.
+After scaling, the three curves are identical in the classical regime (B<0.5B<0.5\,T) as well as on the plateaus of the QHE.
 The line linking shaded and white backgrounds corresponds to the prediction of Eq.
 (II.2) for constant $\mu=\mu_{0}$ [equal to the dashed black lines in Fig.
-7(a)].
+9(a)].
 Compared to the model prediction, between the plateaus the current is reduced for all three geometries, which above we already interpreted in terms of a finite contribution of the non-chiral $I_{xx}$.
-Because the various contact resistances of our sample are almost identical, for a purely chiral current flow the branches of the largest current $I_{k}(B)/I_{1}$, shown in Fig.
-8 should be independent of the details of the contacts configuration (as long as the overall number of grounded contacts remains the same).
-The fact that $I_{k}(B)/I_{1}$ varies between configurations, is a direct experimental prove of non-chiral currents between the quantized plateaus.
-This conclusion experts additional support by a closer look at the amounts of current reduction: It systematically depends on the relative distances between contact 1 and the grounded contacts on both sample edges in a way that contradicts chiral current flow.
+Because the various contact resistances of our sample are almost identical and the number of grounded contacts is identical, for a purely chiral current flow the branches of the largest current $I_{k}(B)/I_{1}$ should be independent of the details of the contacts configuration.
+While this is the case on the quantized Hall plateaus, away from the plateuas $I_{k}(B)/I_{1}$ varies substantially between configurations.
+This can only be understood, if we allow for non-chiral currents between the quantized Hall plateaus.
+This conclusion experts additional support by a closer look at the amounts of current reduction, based on its systematic dependence on the relative distances between contact 1 and the grounded contacts on both sample edges in a way that contradicts chiral current flow.
+Namely, the current supression into the contact with the largest current is the greater the closer another grounded contact is to the biased contact 1.
 III.6 Formulation of the proof of chiral current flow for quantized Hall resistance
 So far, we have proven non-chiral current flow between the quantized plateaus, while for the quantized plateaus we have merely stated that our experimental results are consistent with chiral current flow.
-However, our analysis allows a stronger statement, namely that the agreement between the chiral model and our measurements is a prove of chiral current flow.
+However, our analysis allows a stronger statement, namely that the agreement between the chiral model and our measurements is a prove of chiral current flow, if we accept that Eq.
+(3) describes chiral current flow.
 The proof is based on the coherent and chiral model introduced in Sec.
 II.1.
-Based on the Landauer-Büttiker formalism it describes the current division into $n$ contacts in terms of the mutual conductance coefficients $G_{ij}$ between the contacts.
+Based on the Landauer-Büttiker formalism it describes the current division into $n$ contacts in terms of the mutual conductance coefficients $G_{kj}$ between the contacts.
 In this model, chiral current flow corresponds to the specific values of conductance coefficients given by Eq.
@@
 In all our measurements including three or four terminals shown in Fig.
-5, Fig.
+7, Fig.
 9 and Fig.
-8, we instead see a perfect agreement between our chiral model prediction and the measured plateau values of the current ratios.
-This is in particular evident when comparing the horizontal dotted lines in Fig.
-5(b) and Fig.
-8, which indicate the predictions of Eq.
+10, we instead see a perfect agreement between our chiral model prediction and the measured plateau values of the current ratios.
+This is, in particular, evident when comparing the horizontal dotted lines in Fig.
+7(b) and Fig.
+10, which indicate the predictions of Eq.
 (4) for the corresponding integer filling factors, with the corresponding measured plateau values.
-This result stands in contrast to our finding in between plateaus, where the deviations of our measured currents from the prediction of Eq.
-(II.2) for $\mu B\gg 1$ provide a proof of non-chiral currents.
 IV Summary and Outlook
@@
 We find that the current flow is chiral in the magnetic field regions of the quantized Hall resistance plateaus but becomes non-chiral in between the plateaus.
-Our result is based on the direct comparison of the measured current distribution into various contacts (three or four terminals) with two limiting models, which offer well understood, unique physical meanings.
+Our result is based on the direct comparison of the measured current distribution into various contacts (three or four terminals) with two limiting models, each providing a unique physical meaning.
 The classical Drude model serves us as the basis to describe non-chiral diffusive transport, while our second model is based on the generic chiral and coherent transmission problem.
@@
 Acknowledgement
-The authors thank Piet Brouwer for help with the coherent model and Afif Siddiki for fruitful discussions.
+The authors thank Piet Brouwer for help with the coherent model and Afif Siddiki and Lutz Schrottke for fruitful discussions.
 This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 218453298.
@@
 V. Y. U. supplied the wafer. S. S. and M. K. performed the measurements. S. S. and S. L. analyzed the data. S. L. derived the models. S. S. and S. L. wrote the article.
+Appendix A Currents predicted by our coherent and chiral model
+To derive Eq. (4) in Sec. II.1 describing chiral and coherent current within the inner circle of Fig. 2, rewrite Eq. (2),
+$I_{k}=\sum_{j=1}^{n}G_{kj}V_{j}^{\prime}=(V_{k}^{\prime}-V_{k})/R_{k}\,.$ (11)
+Its left hand side describes the coherent current flowing inside the inner circle of Fig.
+2 according to the Landauer-Büttiker formalism.
+Its right hand side describes the interfaces between the coherent inner circle with the non-coherent world by relating the same currents with the contact resistances based on Ohm’s law.
+Requiring unitarity of the coherent scattering matrix demands $\sum_{k}G_{kj}=\sum_{j}G_{kj}=0$.
+Additionally requiring chiral current flow, such that current flows only from contact $1\to 2$, $2\to 3$, …, $\to n, n1n\to 1$ yields our conductance matrix [equivalent to Eq.
+(3)]
+$\hat{G}=\nu G_{0}\begin{pmatrix}1&0&\cdots&0&-1\\ -1&1&0&\cdots&0\\ 0&-1&1&0&\cdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&-1&1\\ \end{pmatrix}\,$, (12)
+where we used $\nu G_{0}$ for the conductance between pairs of contacts.
+Within the Landauer-Büttiker formalism, $\nu$ stands for the number of coherent channels, each contributing with the one-dimensional conductance quantum, $G_{0}=e^{2}/h$, to the current.
+For the case of the QHE, $\nu$ corresponds to the filling factor and the Hall resistance is $R_{\text{H}}=1/\nu G_{0}$.
+In our experiments, the voltages, $V_{k}$, are applied to the contacts, while we measure the currents, $I_{k}$, flowing into each contact.
+In order to predict these currents, we first determine the voltages, $V^{\prime}_{k}$, at the nodes between the coherent inner circle of Fig.
+2 and the ohmic contacts.
+We can rewrite the right hand side of Eq.
+(3) as
+$\sum_{j=1}^{n}g_{kj}V^{\prime}_{j}=-\frac{V_{k}}{\nu G_{0}R_{k}}\equiv-w_{k}\,$, (13)
+where we introduced the dimensionless conductance matrix
+$\hat{g}=\begin{pmatrix}-\gamma_{1}&0&\cdots&0&1\\ 1&-\gamma_{2}&0&\cdots&0\\ 0&1&-\gamma_{3}&0&\cdots\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&-\gamma_{n}\\ \end{pmatrix}$ (14)
+whith $\gamma_{k}=1+(\nu G_{0}R_{k})^{-1}$. Sorting Eq. (13) for $V^{\prime}_{k}$ results in a unique solution of the linear equation system
+$V^{\prime}_{1}$ $=\frac{1}{\gamma_{1}}\left(V^{\prime}_{n}+w_{1}\right)$
+$V^{\prime}_{2}$ $=\frac{1}{\gamma_{2}}\left(V^{\prime}_{1}+w_{2}\right)$ $=\left(V^{\prime}_{n}+P_{0}w_{1}+P_{1}w_{2}\right)/P_{2}$
+$V^{\prime}_{3}$ $=\frac{1}{\gamma_{3}}\left(V^{\prime}_{2}+w_{3}\right)$ $=\left(V^{\prime}_{n}+P_{0}w_{1}+P_{1}w_{2}+P_{2}w_{3}\right)/P_{3}$
+$\dots$ $\dots$
+$V^{\prime}_{n}$ $=\frac{1}{\gamma_{n}}\left(V^{\prime}_{n-1}+w_{n}\right)$ $=\sum_{j=1}^{n}P_{j-1}w_{j}/\left(P_{n}-1\right)\,$,
+where we introduced the product
+$P_{k}=\left\{\begin{matrix}\prod_{j=1}^{k}\gamma_{j}&\text{for}&j=1,2,3,\dots\\ 1&\text{for}&j=0\end{matrix}\right.\,$. (15)
+With the additional abbreviation
+$S_{k}=\sum_{j=1}^{k}P_{j-1}w_{j}$ (16)
+we find
+$V^{\prime}_{k}=\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)\,.$ (17)
+In the last step, we use $I_{k}=\left(V^{\prime}_{k}-V_{k}\right)/R_{k}$ to express the current into the $k$th contact in terms of the contact resistances $R_{k}$ and the applied voltages $V_{k}$:
+$I_{k}=\frac{1}{R_{k}}\left[\frac{1}{P_{k}}\left(\frac{S_{n}}{P_{n}-1}+S_{k}\right)-V_{k}\right]\,$, (18)
+which is Eq. (4).
+Appendix B Diffusive currents predicted by the Drude model
+Integration over the sample width of the current density in Eq. (II.2) results in
+$I_{k}$ $=$ $I_{xx}^{(k)}+I_{xy}^{(k)}$ (19)
+$\equiv$ $\frac{1}{1+(\mu B)^{2}}I_{x}^{(k)}+\frac{(\mu B)^{2}}{1+(\mu B)^{2}}I_{y}^{(k)}$ (20)
+For determining the ohmic currents $I_{x}^{(k)}$, we assume $V^{\prime}_{1}=V^{\prime}_{2}=\dots=V^{\prime}_{n}\equiv V^{\prime}$ in Fig.
+2, which corresponds to the neglect of the resistance $R_{0}$ of the 2DES.
+Our simplified circuit diagram takes the form of that in cf.
+Fig.
+11.
+Figure 11: Simplified circuit diagram of an $n$-terminal device at $B=0$ neglecting the resistance $R_{0}$ of the 2DES. It describes the current contribution $I_{xx}$ of the Drude model.
+This simplifying ohmic circuit is described by the $n$ equations
+$I_{k}=\left(V^{\prime}-V_{k}\right)/R_{k}\text{ for }k=1,2,\dots,n$ (21)
+and Kirchhoffs current law $\sum_{j=1}^{n}I_{j}=0$. Combining both, we find
+$V^{\prime}=\frac{\sum_{j=1}^{n}V_{j}/R_{j}}{\sum_{j=1}^{n}1/R_{j}}$ (22)
+and
+$I_{x}^{(k)}=\frac{1}{R_{k}}\left(\frac{\sum_{j=1}^{n}V_{j}/R_{j}}{\sum_{j=1}^{n}1/R_{j}}-V_{k}\right)\,$, (23)
+[…]
+Figure 12: Simplified circuit diagram of an $n$-terminal device including the Hall resistance to describe the current contribution $I_{xy}$ of the Drude model.
+The inner circle indicates directional but not chiral current flow through the 2DES from contact to contact.
+Each contact k=1,,nk=1,...,n has an ohmic resistance $R_{k}$ and a voltage $V_{k}$ applied in respect to the electrical ground.
+$V_{k}^{\prime}$ are the voltages resulting at the intersections between the ohmic contacts and the coherent inner circle.
+The magnetic field vector pointing into the plane of the 2DES gives rise to a right handed chirality of a free charge which translates into a left handed chirality of the device, cf.
+circled arrow.
+The currents $I_{y}^{(k)}$ are bent from contact to contact by the charged sample edges as indicated in Fig.
+12.
+Thereby, according to Eq.
+(II.2), inside of the 2DES between contacts, the current $I_{xy}$ is additionally impeded by the Hall resistance $R_{\text{H}}$.
+According to the Drude model, the current contribution $I_{xy}$ flows unidirectional and, thus, is not chiral.
+However, it flows only between adjacent contacts in a pre-defined direction.
+The model describing $I_{xy}^{(k)}$ is mathematically equivalent with our coherent and chiral model, if we use $\nu G_{0}=1/R_{\text{H}}$.
+However, while the coherent model applies for integer filling factors only, the pre-conditions of the Drude model exclude the limit of integer $\nu$.
+Appendix C Hall measurements
+In Fig. 13,
+Figure 13: Standard Hall measurements, $R_{ij}=V_{ij}/I_{41}$ of our Hall bar at the temperature of $T\simeq 300\,$mK, current applied between contacts 1 and 4, cf.
+inset.
+Hall resistances, $R_{\text{H}}(B)$, in the top panel and longitudinal resistances, $R_{\text{x}}(B)$, in the bottom panel.
+[…]
+The slight deviations between the longitudinal resistances, $R_{23}$ and $R_{65}$, in Fig.
+13 are not reproducable and might be related with the fact that we measured the voltages at the four contacts with four separate volt meters.
+A slight shift in the offset of one of them between measurements might cause the observed differences.